### v2

# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

Abstract
Verifying the provenance of content is crucial to the function of many organizations, e.g., educational institutions, social media platforms, firms, etc. This problem is becoming increasingly difficult as text generated by Large Language Models (LLMs) becomes almost indistinguishable from human-generated content. In addition, many institutions utilize in-house LLMs and want to ensure that external, non-sanctioned LLMs do not produce content within the institution. In this paper, we answer the following question: Given a piece of text, can we identify whether it was produced by LLM $A$ or $B$ (where $B$ can be a human)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design zero-shot statistical tests to distinguish between (i) the text generated by two different sets of LLMs $\mathcal{A}$ (in-house) and $\mathcal{B}$ (non-sanctioned), and also (ii) LLM-generated and human-generated texts. We prove that the type I and type II errors for our tests decrease exponentially in the text length. In designing our tests, we derive concentration inequalities on the difference between log-perplexity and the average entropy of the string under $A$. Specifically, for a given string, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. We also show that if $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Lastly, we present preliminary experimental results using open-source LLMs to support our theoretical results. Practically, our work enables guaranteed (with high probability) finding of the origin of harmful or false LLM-generated text for a text of arbitrary size, which can be useful for combating misinformation as well as compliance with emerging AI regulations.
1 Introduction
LLM and human-generated texts are becoming indistinguishable. This phenomenon has some concerning societal consequences, including the spread of LLM-generated misinformation and LLM-assisted academic cheating. We need reliable detection methods to distinguish between human and AI-generated content. In addition to distinguishing human vs. LLM-generated text, the ability to differentiate between text generated by a language model vs. another language model is also critical because of the following reasons. First, the ability to find the origin of harmful or false LLM-generated content is essential for legal compliance and mitigation purposes (Hacker et al., 2023; Wang et al., 2023). Identifying the source of harmful content allows for responsibility to be assigned in case of non-compliance with regulations. Second, the quality of outputs can vary significantly between models, and as a result, employing the most appropriate model for the specific task is essential to achieve optimal results. In line with operating the most suitable language model, educational organizations are building their LLMs to secure educational integrity and credibility. For example, the University of Michigan has developed U-M GPT and UM Maizey as its generative AI tools to ensure academic integrity, guarantee user data protection, and assure that the shared information does not train the underlying AI models (UMichigan (2023)). Students are allowed only to use the specialized assistant (here, UM Maizey) to do their assignments, not ChatGPT or any other LLM. This requirement highlights the need for reliable tools to detect the text generated by prohibited LLMs.
To evaluate the possibility of detecting LLM-generated text by human experts in linguistics, Casal and Kessler, 2023 designed an experiment to investigate whether linguists can distinguish human and ChatGPT-generated text and reported an identification rate of only $38.9\%$. Since humans, even experts, perform poorly in detecting LLM-generated text, researchers are investing significant efforts in designing automated detection methods to identify signals that are difficult for humans to recognize.
One way to create a detection method is to train classifiers on labeled training data coming from LLM-generated and human-generated classes. OpenAI itself released an AI classifier model. However, they later discontinued operating the model due to its low accuracy rate (OpenAI ). While recent literature on classification methods with higher accuracy rates exists (e.g., Guo et al., 2023), the limitations of this method make its application practically challenging. The first major limitation is the requirement for training a separate (from the source-model) classifier, especially considering the large and growing number of LLMs, the wide variety of topics and writing styles, and the possibility of prompting LLMs to write in different styles. Furthermore, the requirement for collecting a dataset of human and AI generated passages raises concerns, such as privacy, associated with training models on human data. Finally, Liang et al., 2023 note that because detectors are often evaluated on relatively easy datasets, their performance on out-of-domain samples is often abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using commercial detectors.
An alternative approach for detecting LLM-generated content is watermarking. However, watermarking relies on cooperation from the AI company/owner of the LLM. Although using watermarking for AI regulation is imaginable, current regulations cannot force companies to adopt this technology (Nature, 2024). The described limitations motivate the need for models that do not require training on human data or cooperation from the LLM owner. One such method is zero-shot training (see, e.g., the work of Mitchell et al., 2023).
The majority of the zero-shot detection studies (e.g., Mitchell et al., 2023 and Hans et al., 2024) leverage statistical properties of LLM-generated texts to develop statistical tests for identifying whether a human or an LLM generates a finite-length text. The proposed methods rely on heuristics that perform well in practice, but they do not provide statistical guarantees for their detection mechanism’s accuracy level, making them unfit for any purpose where strong evidence is required (e.g., regulatory compliance).
In this study, we answer the following question: Given a piece of text, can we identify whether it was generated by LLM $A$ or $B$ (where $B$ can be a human or another LLM among a set of LLMs)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single string of text with finite length, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We design tests to distinguish between different LLMs. For this purpose, we assume that we have white-box access to the models in the hypothesis test. In particular, we design composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. We also study the case where we do not have white-box access to all models in the hypothesis set (for example, a human wrote the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
We contribute to the literature on zero-shot statistical tests by developing the first statistical test (with theoretical guarantees) that identifies whether a finite-length text was generated by an LLM or by a human. We show that the type I and type II errors for our statistical tests decrease exponentially as the text length increases.
With the development of specialized LLMs such as UM Maizey, enabling a theoretically guaranteed linking of a finite-length text to its origin among a set of LLMs is becoming necessary. We are the first to provide statistical tests with guarantees for this problem, and we also prove that the type I and type II errors for our statistical test decrease exponentially as the text length increases.
Finally, our theoretical results include establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables. These concentration bounds have been derived for a sequence of independent and identically distributed (iid) random variables by Zhao 2022. However, large language models generate text sequences dependent on the previously generated tokens. Motivated by the problem requirement, we generalize the results in the literature (e.g., Zhao 2022) by proving an exponential decay concentration inequality to bound the tail probability of the difference between the log-likelihood of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables satisfying a martingale structure. Interestingly, as a byproduct of this paper, we address one of the future research directions mentioned by Zhao 2022, namely deriving entropy concentration bounds for martingales.
This paper is organized as follows: In Section 2, we present our mathematical framework and define the critical random variables necessary for deriving our theoretical results. Section 3 consists of two subsections. In Subsection 3.1, we provide concentration bounds when the generator and evaluator models are the same, and in Subsection 3.2, we provide concentration bounds when the generator and evaluator models are different. In Section 4, we present our statistical tests and derive the upper bound on their type I and type II errors. In Section 5, we show the results of our preliminary experiments on an open-source language model. We conclude the paper in Section 6. Appendix A presents the detailed proofs, and Appendix B provides a literature review.
2 Model and background definitions
2.1 Model
Let $M$ be a generative model described by $\textbf{Y}=m(\textbf{X})$, where X denotes user prompt and the output denoted by Y consists of a string of tokens $\textbf{Y}=[Y_{1},Y_{2},\dots,Y_{N},\dots]$. Each token is chosen from a finite vocabulary set, i.e., $Y_{n}\in\mathcal{X}$, and we denote the vocabulary size by $K:=|\mathcal{X}|$.
Practical implementations of LLMs specify the probability distribution iteratively, e.g., Radford 2018. The model first draws a random value for the first token, say $Y_{1}=y_{1}$ by sampling from the distribution $p^{M}(Y_{1}|\textbf{X})$, and then for each token $n\in[2,N]$, the model sequentially determines a distribution for the token given prompt X and all the randomly chosen values $y_{1},y_{2},\dots,y_{n-1}$. So, we define a sequence of probability distributions $p^{M}(\textbf{Y}_{N}|\textbf{X})$ over $\textbf{Y}_{N}\in\mathcal{X}^{N}$ where $\textbf{Y}_{N}=[Y_{1},Y_{2},\dots,Y_{N}]$ is a substring of Y consisting of the first $N$ tokens. The sequence of probability distributions is determined as
$P^{M}(\textbf{Y}_{N}|\textbf{X})=\prod_{n=1}^{N}p^{M}_{n}(\textbf{Y}_{n}),\text{where}\quad p^{M}_{n}(\textbf{Y}_{n})=P^{M}(Y_{n}|Y_{1},Y_{2},\dots,Y_{n-1},\textbf{X}).$ (1)
Remark 1.
Equation (1) is an application of the Bayes’ rule and holds for any generative model (regardless of whether tokens $Y_{n}$ are sequentially generated). While equation (1) holds for all generative models, because conditional distributions $p_{n}(y)$ are in general not easily accessible, we apply the rule for sequential models. We follow the literature on white-box detection, and assume that we have complete knowledge of the probability law $p^{M}_{n}(\textbf{Y}_{n})$ for any given sequence $\textbf{Y}_{n}$. See, for example, Mitchell et al. 2023; Gehrmann et al. 2019.
2.2 Background definitions
The perplexity $p^{A}(Y_{N})$ of a (finite length) text string $\textbf{Y}_{N}=[Y_{1},Y_{2},\ldots,Y_{N}]$ with respect to an evaluator generative model $A$ is defined as the per-token inverse likelihood of the string Y. Formally, perplexity with respect to model $A$ is
$p^{A}(\textbf{Y}_{N})=\left(\prod_{n=1}^{N}p^{A}_{n}(Y_{n})\right)^{-\frac{1}{N}},$
and log-perplexity with respect to model $A$ is
$l_{A}(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\log(p^{A}_{n}(y_{n})).$ (2)
Intuitively, log-perplexity measures how "surprising" a string is to a language model.
We aim to design statistical tests to reliably evaluate whether text is generated by model $A$ or by a different model $B$, which can be a different generative model or human. Suppose model $B$ generates a string $\mathbf{Y}$. The cross-entropy of model $B$ and evaluator model $A$ over sub-string $\mathbf{Y}_{N}$ equals
$h_{N}(B,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$ (3)
Note that if model $A$, which is also the evaluator model, generates the string, then cross-entropy and entropy are equivalent and defined as
$h_{N}(A,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}p^{A}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$
Our consideration of white-box detection yields that we have complete knowledge of the probabilities $p^{A}_{n}(y_{n})$, and therefore, we can compute $h_{N}(A,A)$. Furthermore, Gibbs’ inequality (Cover (1999)) states that $h_{N}(B,A)\geq h_{N}(B,B)$.
2.3 Key random variables
Here, we define the random variables that are critical in deriving our theoretical results.
Let us first define the random variable $Z_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$. If $\mathbb{E}[Z_{n}]<\infty$, then we define a zero-mean random variable $X_{n}=:Z_{n}-\mathbb{E}[Z_{n}]$. If model $A$ generated the string, then we denote the expected value for the random variable $Z_{n}$ by $\mathbb{E}_{p_{n}^{A}}[Z_{n}]$. Lastly, define a random variable $S_{N}:=\sum_{i=1}^{N}X_{i}$.
3 Concentration bounds
In this section, we present our results in two parts. In Section 3.1, we provide concentration bounds to show that if the string is generated by model $A$, $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average entropy of the string under $A$ with a high probability. In Section 3.2, we provide concentration bounds to show that if the string is generated by another model $B$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average cross-entropy of the string under $B$ and $A$ with a high probability. These concentration bounds are the backbones of the statistical tests that we design in section 4.
3.1 Same generative and evaluator models
Consider a string $\mathbf{Y}$ generated by model $A$ and we evaluate the text using the same model $A$. First recall that given a string $\mathbf{Y}_{N}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$. Then,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n}).$
Also, given the string $\mathbf{Y}_{N}$, since we have complete knowledge on the probabilities $p^{A}_{n}(y_{n})$, we have complete knowledge on the entropy $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})=\mathbb{E}_{p_{n}^{A}}[Z_{n}]$. Then, we can find the following upper bound for $\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Remark 2.
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})\leq\log|\mathcal{X}|=\log(K).$
Proof.
By concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, the value of its maximizer is $p^{*A}_{n}(y_{n})=\frac{1}{|\chi|},\quad\forall y_{n}\in\chi$. Thus,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log p_{n}^{A}(y_{n})$$\leq$ $\sum_{y_{n}\in\chi}p_{n}^{*A}(y_{n})\log p_{n}^{*A}(y_{n})$
$=$ $|\chi|.\frac{1}{|\chi|}.\log|\chi|=\log(K).$
With $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Lemma 1.
The random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale.
Proof.
$Z_{n}$ and as a result $\mathbb{E}_{p_{n}^{A}}(y_{n})$ are positive random variables, and we showed that $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\log(K)$. Therefore, we have $\mathbb{E}\big{[}|X_{n}|\big{]}\leq\log(K)$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}|S_{N}]=S_{N}$.
(i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}=\mathbb{E}\big{[}|X_{N+1}|\big{]}\leq\log(K)<\infty$.
(ii) is satisfied because the martingale increments $X_{n}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Finally, we apply concentration bounds for martingales to provide finite sample guarantee for the convergence of the random variable $S_{N}/N$ to zero. A challenge in applying the common concentration bounds for martingales is that martingale increments are not necessarily bounded. We overcome this issue by showing that the martingale differences, while not bounded, admit a light tail. In particular, we show that the martingale differences are sub-exponential.
Definition 1.
(sub-exponential norm). The sub-exponential norm of $X\in\mathbb{R}$ is
$\|X\|_{\psi_{1}}=\inf\bigg{\{}t>0:\mathbb{E}[e^{\frac{|X|}{t}}]\leq 2\bigg{\}}.$
If $\|X\|_{\psi_{1}}$ is finite, we say that $X$ is sub-exponential.
Lemma 2.
The sub-exponential norm for random variable $X_{n}$ equals $2\log(K)$.
Proof.
See Appendix A.1. ∎
As the last step, we apply the concentration bounds for martingales with sub-exponential increments to obtain the following concentration bound.
Theorem 1.
There exists a constant $c_{1}>0$ independent of the evaluator model $A$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}$
Proof.
See Section A.2. ∎
Interpretation. Theorem 1 states that if a given string is generated by a model same as the evaluator (here, model $A$), then the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length.
3.2 Different generative and evaluator models
In this section, we consider a string $\bf{Y}$ generated by model $B$ and want to evaluate our statistical test based on model $A$. To design the statistical test, first recall that given a string $\bf{Y}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(y_{n})\big{)}$. Note that
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))=H(p^{B}_{n},p^{A}_{n}),$ (4)
where $H(p^{B}_{n},p^{A}_{n})$ is the cross entropy between the two distributions $p^{B}_{n}(.)$ and $p^{A}_{n}(.)$.
Note that, unlike the case analyzed in section 3.1, here, $\mathbb{E}_{p_{n}^{B}}[Z_{n}]$ is not necessarily finite.
For $\mathbb{E}_{p^{B}_{n}}[Z_{n}]$ to be infinite, as we can infer from equation 4, we must have that $p^{A}_{n}(y_{n})=0$ and $p^{B}_{n}(y_{n})>0$ for some $y_{n}\in\chi$. In this case, if the string includes such $y_{n}$, then we realize that the string is not generated by model $A$ with the probability of 1. This is a trivial case.
Yet, if the string does not include any such $y_{n}$, then we can update the probability distributions as
$\tilde{p}^{B}_{n}=\frac{p^{B}_{n}(y_{n})}{\sum_{y_{k}:p^{A}_{n}(y_{k})>0}p^{B}_{n}(y_{k})}.$
Without loss of generality, we can exclude the trivial case and assume that if $p^{B}_{n}(y_{n})>0$, then $p^{A}_{n}(y_{n})>0$, which yields that $\mathbb{E}_{p^{B}_{n}}[Z_{n}]=H(p^{B}_{n},p^{A}_{n})$ is finite. With $\mathbb{E}_{p_{n}^{B}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p^{B}_{n}}[Z_{n}]$. For tractability, we make a parametric assumption on the probability laws $p^{A}_{k}(.)$ and $p^{B}_{k}(.)$.
Assumption 1.
For our two model analysis, we assume that there exists $\epsilon>0$ such that $p^{A}_{n}(y_{k}),p^{B}_{n}(y_{k})\notin(0,\epsilon)$.
Assumption 1 implies that models $A$ and $B$ either do not associate any probability to a token $y\in\mathcal{X}$, or they assign a probability of at least $\epsilon$. Our theoretical results depend only on $\log(\epsilon)$. Hence, our theoretical bounds primarily rely on a constant shift in the logarithmic scale.
It is noteworthy that Assumption 1 is not restrictive and is aligned with practice, as computers only allow for a limited range of representable numbers due to finite precision in floating-point arithmetic. Very small probabilities are either rounded to zero or set to a minimum threshold to maintain numerical stability in computations (Goldberg, 1991).
The first outcome of assumption 1 is that
$\mathbb{E}_{p^{B}_{n}}\big{[}|Z_{n}|\big{]}=\mathbb{E}_{p^{B}_{n}}\big{[}Z_{n}\big{]}=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon)$ (5)
Hence, with the same argument as in section 3.1, the random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale. To apply the martingale concentration bounds, similar to the previous section, we first find the sub-exponential norm for the random variable $X_{n}$.
Lemma 3.
Under Assumption 1, the sub-exponential norm for the random variable $X_{n}$ equals $-4\log(\epsilon)$.
Proof.
See Section A.3. ∎
Theorem 2.
There exists a constant $c_{3}>0$ independent of models $A$ and $B$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}$
Proof.
See Section A.4. ∎
Interpretation. Theorem 2 states that if model $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under model $A$ converges to the average cross-entropy of the string under $B$ and $A$.
4 Statistical test
Now we design our statistical tests using the results in Theorems 1 and 2 and then evaluate type I (false positive) and type II (false negative) errors. In particular, we consider a finite-length text with length $N$ generated by a model $M$. We first design tests for detection between different LLMs. In that, we have white-box access to all models in the hypothesis test. In Section 4.1, we design a simple statistical test that determines whether a text is generated by a model $A$ or another model $B$. Then, in Section 4.2, we extend our results to composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. Finally, in Section 4.3, we study the case where we don’t have white-box access to all models in the hypothesis set (for example, a human produced the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
4.1 Simple statistical test for detection between two LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with length $N$, we design a statistical test to detect whether model $A$ or model $B$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by $B$, and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by $A$. We first calculate the random variables $Z^{A}_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$ and $Z^{B}_{n}=:-\log\big{(}p^{B}_{n}(Y_{n})\big{)}$, and then we calculate the sums $\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}$, and $\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the model $A$ and incorrectly concludes that it is generated by $B$.
To quantify our model’s type I and type II errors, we need to make the following (mild) assumption.
Assumption 2.
(minimum difference). We assume that if the generative and evaluator models are different, for an arbitrarily small positive $\epsilon_{1}>0$, we have
$\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\geq\epsilon_{1}.$
Assumption 2 ensures that the two models satisfy a minimum distance in terms of their KL divergence over the generated text. Note that KL divergence, by definition, is a non-negative value that demonstrates the distance between the two distributions over the next word for the two models. Our results show that the type I and type II errors of our statistical test are approximately $\exp\big{(}O(-N\epsilon_{1})\big{)}$, which indicates that even for small values of $\epsilon_{1}$ that can converge to zero with the length of text (for example, $\epsilon_{1}=O(N^{-1/2})$), our statistical test provides exponentially small type I and type II errors in the length of the text. Hence, our theoretical bounds only require that the two models do not impose the same probability distribution on the string.
Proposition 1.
If Assumptions 1 and 2 hold, then the type I and type II errors for our statistical test are upper bounded by
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.5. ∎
Interpretation. Proposition 1 demonstrates that the type I and type II errors of our simple test decrease exponentially in the text length.
4.2 Statitical test for detection among multiple LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, we design a statistical test to detect whether the text is generated by one of the models $\mathcal{A}=\{A_{1},\ldots,A_{p}\}$ or one of models =$\mathcal{B}=\{B_{1},\ldots,B_{q}\}$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by one of the models in $\mathcal{B}$, and the alternative hypothesis $\mathbf{H}_{1}$ is that it is generated by one of the models in $\mathcal{A}$. We first calculate the random variables $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, and sum $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, for all models $M\in\mathcal{A}\cup\mathcal{B}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if for some $A_{i}\in\mathcal{A}$, we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n};\qquad\forall B_{j}\in\mathcal{B}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by one of the models in $\mathcal{A}$ when it is written by one of the models in $\mathcal{B}$, and Type II error happens when the test fails to identify that text is generated by one of the models in $\mathcal{A}$ and incorrectly concludes that it is generated by one of the models in $\mathcal{B}$. Similar to our test for the two model version, Assumption 2 must hold for us to quantify our model’s type I and type II errors.
Proposition 2.
If Assumptions 1 and 2 hold, the type I error for our statistical test is upper bounded by
$2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
and the type II error for our statistical test is upper bounded by
$2|\mathcal{B}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.6. ∎
Interpretation. Proposition 2 demonstrates that the type I and type II errors of our composite test decrease exponentially in the text length.
4.3 Statistical test for detection between an LLM and human
We need to make the following assumption to design our test.
Assumption 3.
(minimum tangible difference). We assume that if the generative and evaluator models are different, then
$\frac{1}{N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}\textbf{Y}_{n-2}\bigg{]}\geq 4\log^{2}\big{(}K\big{)}.$
Assumption 3 ensures that the two models satisfy a minimum distance in terms of their expected KL divergence. Clearly, if models are the same (or very similar), then $A$ and $B$ impose the same (or almost the same) probability distributions $p_{N}^{A}(y_{n})$ over $\mathbf{Y}_{N}$, and hence KL divergence becomes zero. This makes differentiation impossible. Assumption 3 rules out cases where the two models impose very similar distributions over the text under evaluation.
Lemma 4.
Under assumption 3, for any positive constant $c_{4}\leq\log^{2}(K)/2$, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Proof.
See Section A.7. ∎
Remark 3.
We note that the only application of Assumption 3 for establishing our results is that this assumption ensures that the statement $c_{4}\leq\log^{2}(K)/2$ in Lemma 4 holds. Our results remain true ( even without Assumption 3) if there exists $c_{4}>0$ such that if the generative and evaluator models are different, then
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, for arbitrary constants $t<c_{4}\leq\log^{2}(K)/2$, we design a statistical test to detect whether the evaluator model $A$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is not generated by the evaluator model $A$ (e.g., it is generated by another model $B$), and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by the evaluator model $A$. We first calculate the random variable $Z_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and then we calculate the sum $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ in favor of the alternative $\mathbf{H}_{1}$ if
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the evaluator model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the evaluator model $A$ and incorrectly concludes that it is not written by $A$.
Proposition 3.
For any $t\geq 0$, the type II error for our statitical test is upper bounded by
$2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
with $c_{1}$ as introduced in Theorem 1.
Additionally, if Assumptions 1 and 3 hold, then for any positive $t<c_{4}$, type I error for our statistical test is upper bounded by
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.,$
with $c_{3}$ as introduced in Theorem 2 and $c_{4}$ as introduced in Assumption 3.
Proof.
See Section A.8. ∎
Interpretation. Proposition 3 demonstrates that the type I and type II errors of our test for detecting LLM $A$ vs. not LLM $A$ (that includes human) decrease exponentially in the text length.
5 Preliminary experiments
We conduct our numerical analysis on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
5.1 Same generative and evaluator model
In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (1(a)-1(d)). We consistently observe that the numerical results confirm Theorem 1 that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
(a) GPT-2 small
(b) GPT-2 medium
(c) GPT-2 large
(d) GPT-2 XL
Figure 1: Generated and evaluated by the same model
5.2 Different generative and evaluator models
To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (2(a)-2(c)). Results in these figures confirm Theorem 2. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XL
Figure 2: Generated by GPT-2 small and evaluated by different models
6 Conclusion
In this study, we establish first zero-shot statistical tests with theoretical guarantees for text with finite length to distinguish between (i) LLM-generated and human-generated texts in Proposition 3, and (ii) the text generated by two different LLMs $A$ and $B$ in Propositions 1 and 2. We prove that the type I and type II errors for our tests decrease exponentially in the text length. As a critical step in designing our tests, we derive concentration bounds in the difference between log-perplexity and the average entropy of the string under $A$. Specifically, for a given string, in Theorem 1, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. Furthermore, in Theorem 2, we show that if $B$ (which can be either another model or human) generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Our theoretical results rely on establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables and the negative entropy for non-independent random variables on a finite alphabet. Results in the literature (e.g., Zhao (2022)) derive concentration bounds for iid random variables, and one of our theoretical contributions is to extend the results to non-independent random variables by introducing random variables that form a martingale. We hope that our work inspires more research on zero-shot LLM-text detection with provable guarantees.
Appendix A Appendix
A.1 Proof of Lemma 2
Proof.
For the random variable to be sub-exponential, by definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\frac{1}{p^{A}_{n}(y_{n})^{1/t}}=e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq 2.$ (6)
where the first inequality follows from the triangle inequality and that $Z_{n}$ and therefore $\mathbb{E}[Z_{n}]$ are positive. Note that $\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}$ is concave in $p^{A}_{n}(y_{n})$. Hence, it attains its maximum when we have $p^{A}_{n}(y_{i})=p^{A}_{n}(y_{j})\quad\forall y_{i},y_{j}\in\mathcal{X}$, which yields
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq K(\frac{1}{K})^{\frac{t-1}{t}}=K^{\frac{1}{t}}.$ (7)
To analyze $e^{\frac{\mathbb{E}[Z_{n}]}{t}}$, we first want to show that
$\mathbb{E}[Z_{n}]$ $=$ $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ (8)
$\leq$ $\log|\mathcal{X}|=\log(K).$
To show that, note that by concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, we can maximize $p_{n}(y_{n})\log p_{n}(y_{n})$ by equalizing all probabilities $p^{*}(y_{n})=\frac{1}{|\mathcal{X}|};\forall n$. So, we have
$-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$$\leq$ $-\sum_{y_{n}\in\mathcal{X}}p^{*}(y_{n})\log p^{*}(y_{n})$
$=$ $-|\mathcal{X}|\Big{[}\frac{1}{|\mathcal{X}|}\log\frac{1}{|\mathcal{X}|}\Big{]}=\log|\mathcal{X}|.$
Combining equations (6)-(8) yields that for $X_{n}$ to be sub-exponential with norm $t$, it must satisfy
$K^{1/t}K^{1/t}\leq 2$
Hence, $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=2\log(K)$. ∎
A.2 Proof of Theorem 1
Proof.
We first present an equivalent (up to a constant factor) definition of a sub-exponential random variable from Vershynin (2018).
Definition 2.
(sub-exponential random variable). Centered random variable $X\in SE(\nu^{2},\alpha)$ with parameters $\nu,\alpha>0$ is sub-exponential if
$\mathbb{E}[e^{\lambda X}]\leq e^{\frac{\lambda^{2}\nu^{2}}{2}},\quad\forall\lambda:|\lambda|<\frac{1}{\alpha}.$
Next, we present a lemma from Vershynin (2018) that demonstrates that the two definitions are equivalent up to a constant factor.
Lemma 5.
(SE properties ,Vershynin, 2018). Let $X$ be a random variable with $\mathbb{E}[X]=0$. Then, there exists a constant $c$ and constants $K_{4}$ and $K_{5}$ such that $K_{4}\leq cK_{5}$ and $K_{5}\leq cK_{4}$ and the following two properties are equivalent.
  • There exists a constant $K_{4}$, such that the MGF of $|X|$ is bounded, specifically
    $\mathbb{E}[e^{|X|/K_{4}}]\leq 2.$
  • There exists a constant $K_{5}$, such that the MGF of $X$ satisfies
    $\mathbb{E}[e^{\lambda X}]\leq e^{\frac{K_{5}^{2}\lambda^{2}}{2}}\quad\forall\lambda\quad s.t.\quad|\lambda|\leq\frac{1}{K_{5}}$
Next, note that by Lemma 2 and Lemma 5, there exists $c_{1}>0$ s.t. for $\nu=c_{1}\log K$, and for any $n\in\mathbb{N}$,
$\mathbb{E}[e^{\lambda X_{n}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (9)
Let $\alpha=\nu$. By Definition 2, we have $X_{n}\in SE(\nu^{2},\alpha)$.
Our next step is to show that $\sum_{n=1}^{N}X_{n}$ is SE with parameters $(\nu\sqrt{N},\alpha_{*})$. To realize that, observe that
$\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n}X_{k}\right)}\right]=\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n-1}X_{k}\right)}\mathbb{E}\left[e^{\lambda X_{n}}\mid\mathbf{Y}_{n-1}\right]\right]\leq\mathbb{E}\left[e^{\lambda\sum_{k=1}^{n-1}X_{k}}\right]e^{\frac{\lambda^{2}\nu^{2}}{2}}\leq e^{\frac{\lambda^{2}N\nu^{2}}{2}},$
where the first equation follows from the iterated law of expectation, and the first inequality follows from Equation (9).
Finally, from Theorem 5.2 in Arinaldo, 2019, if $S\in SE(\nu^{2},\alpha)$ is a sub-exponential random variable, then
$\mathbb{P}(|S-\mathbb{E}[S]|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{\nu^{2}},\frac{t_{1}}{\alpha}\bigg{)}\bigg{)}.$ (10)
Substituting $\nu=\alpha\sqrt{N}$ for the zero-mean random variable $\sum_{n=1}^{N}X_{n}$, and $\alpha=c_{1}\log K$, we obtain
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(c_{1}\log K)^{2}},\frac{t_{1}}{c_{1}\log K}\bigg{)}\bigg{)}.$
Setting $t_{1}=tN$, we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{i=1}^{N}X_{i}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.3 Proof of Lemma 3
Proof.
For the random variable to be sub-exponential, by definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{p^{A}_{n}(y_{n})^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{\epsilon^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}}\leq 2.$
Thus, for $X_{n}$ to be sub-exponential with norm $t$, it is sufficient to satisfy
$\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}\leq\log(2)\leq 1/2.$ (11)
Recall equation 5 that states
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon).$
Substituting the result of (5) in (11) yields that $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=-4\log(\epsilon)$. ∎
A.4 Proof of Theorem 2
Proof.
Following the same steps as in proof of Theorem 1, we conclude that $\sum_{n=1}^{N}X_{n}$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t_{1}}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Finally, setting $t_{1}=tN$, we obtain
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
A.5 Proof of Proposition 1
Proof.
Type I error occurs if the model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
By triangle inequality, this yields
$h_{N}(B,A)-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq h_{N}(B,B)-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B,A)-h_{N}(B,B)\geq\epsilon_{1}$. Hence, the type I error occurs only if
$\epsilon_{1}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Equivalently, the type I error occurs only if
$\epsilon_{1}\leq\bigg{|}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\bigg{|}+\bigg{|}-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}\bigg{|}.$
So, the type I error only occurs if at least $|-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 2, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
The type II error occurs if the model $A$ generates the text $\bf{Y}_{N}$, but we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}.$
By symmetry, the type II error adheres to the same upper bound as the type I error, yielding
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.6 Proof of Proposition 2
Proof.
Type I error occurs if a model $B_{j}\in\mathcal{B}$ generates the text string $\mathbf{Y}_{N}$, but for one model $A_{i}\in\mathcal{A}$ we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
This yields for some $A_{i}\in\mathcal{A}$, we have
$h_{N}(B_{j},A_{i})-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq h_{N}(B_{j},B_{j})-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B_{j},A_{i})-h_{N}(B_{j},B_{j})\geq\epsilon_{1}$. Hence, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
Equivalently, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}+\bigg{|}-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}.$
So, the type I error only occurs if, for at least one of the models $A_{i}\in\mathcal{A}$, we have either $|-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+|\mathcal{A}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 2, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}+2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if a model $A_{i}\in\mathcal{A}$ generates the text $\bf{Y}_{N}$, but for a model $B_{j}\in\mathcal{B}$ we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
This yields that for some $B_{j}\in\mathcal{B}$, we have
$h_{N}(A_{i},B_{j})-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq h_{N}(A_{i},A_{i})-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Note that by Assumption 2, we have $h_{N}(A_{i},B_{j})-h_{N}(A_{i},A_{i})\geq\epsilon_{1}$. Hence, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Equivalently, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}+\bigg{|}-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}.$
So, the type II error only occurs if at least for one of the models $B_{j}\in\mathcal{B}$, we have either $|-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type II error as
$|\mathcal{B}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 2 we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Also, from Theorem 1 we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Therefore, the type II error is at most
$2|\mathcal{B}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.7 Proof of Lemma 4
Proof.
We prove this lemma in three steps.
Step 1. sub-exponential norm for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
For the random variable $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$, applying (5) we obtain
$D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}=\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\big{(}\log(p^{A}_{n}(y_{n}))-\log(p^{B}_{n}(y_{n}))\big{)}\leq\sum_{y_{n}\in\mathcal{X}}-p^{A}_{n}(y_{n})\log(p^{B}_{n}(y_{n}))\leq-\log(\epsilon),$ (12)
where the first inequality holds since $\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log(p_{n}^{A}(y_{n}))\leq 0$, and the second inequality holds by Assumption 1.
For the random variable to be sub-exponential, by Definition 1, we need to find $t$ such that
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq 2.$
Applying (12), for the random variable to be sub-exponential with norm $t$, it is sufficient to satisfy
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq e^{\frac{-2\log(\epsilon)}{t}}\leq 2.$
Hence, $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ has sub-exponential norm $-4\log(\epsilon)$.
Step 2. Concentration bounds for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
Following the same steps as in the proof for Theorem 1, we conclude that $\sum_{n=1}^{N}[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Next, setting $t=c_{4}N$, we obtain
$\mathbb{P}\bigg{(}\big{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\big{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (13)
Step 3. Tail bound for cross-entropy
In this proof, for notation brevity we write $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$ instead of $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$, and $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\big{]}$ instead of $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}\big{]}$.
First, by Theorem 1 in Reeb and Wolf (2015), we have
$\Big{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}\leq\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K).$ (14)
Therefore, we have
$\Big{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}$ (15)
$=$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\bigg{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K)$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}$
$\geq$ $\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}\bigg{(}\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}-\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}\bigg{)}$
$\geq$ $\frac{1}{2N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{4N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]},$
where the first equality uses the definition of cross-entropy. The first inequality follows from the triangle inequality. The second inequality follows form (14). The third inequality follows from Assumption 3. The fourth inequality follows from QM-AM inequality. The last inequality follows from the difference of squares identity.
Hence, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\leq 2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}-\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq c_{4}\bigg{)}$
$\leq$ $2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]},$
where the first inequality follows from (15). The fourth inequality holds under Assumption 3, which requires $\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq 2\log^{2}(K)$ and also under the assumption in the statement of Lemma 4, which requires that $2c_{4}\leq\log^{2}(K)$. Finally, the last inequality follows from (13). This concludes the proof. ∎
A.8 Proof of Proposition 3
Proof.
Type I error occurs if a model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Applying triangle inequality, we have
$\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}.$
Hence, the type I error is upper bounded as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$ (16)
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}.$
From Lemma 4, we know that
$\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (17)
From Theorem 2 we know that
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (18)
Combining Equations (16-18), we conclude the type I error is upper bounded as
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if the model $A$ generates the text $\bf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t.$
We upper bound the probability of this event as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
where the inequality follows from Theorem 1. ∎
Appendix B Related literature
Preemptive detection methods. One detection approach is to record (Krishna et al., 2024) or watermark (Kirchenbauer et al., 2023) all generated text. The method proposed in Kirchenbauer et al. (2023) takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. For a survey on watermarking methods, we refer readers to Amrit and Singh (2022). It is noteworthy to mention that preemptive approaches can only be applied if the LLM owner cooperates by allowing full control over the generative model.
Unlike preemptive detection methods, post-hoc detection methods can be used without needing cooperation from model developers. The two main categories in this group are neural-based and statistical-based methods.
Neural-based methods. Significant efforts have been made to develop neural-based methods (e.g., features-based classifiers). However, a key concern with this approach is poor robustness, e.g., against ambiguous semantics (Schaaff et al. (2023)) and the neural-based method’s limited ability to detect LLM-generated misinformation (Schuster et al. (2020)). Recent efforts have been enhancing training methodology (e.g., Kumarage et al., 2023; Tu et al., 2023). However, these methods provide only heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts. Additionally, there are limitations for the feature-based classifiers, including challenges in training models with the hype in developing new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns.
Statistical-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. The majority of works in this category also only provide heuristics.
  • White-Box Statistics. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures (Solaiman et al. (2019)). Other measures include using GTLR based on rank-likelihood (Gehrmann et al. (2019)), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. (2023), entropy and Kullback-Leibler (KL) divergence (e.g., Lavergne et al. (2008)), and perplexity (e.g., Verma et al., 2021; Vasilatos et al., 2023). The majority of research in applying white-box statistics proposes heuristics.
  • Black-Box Statistics. Based on the observation that human-generated texts require more revisions when LLMs are tasked with editing the text compared to LLM-generated texts, several works, including Yang et al., 2023; Zhu et al., 2023; Mao et al., 2024 employ LLMs to continue writing truncated texts under review and differentiate between human and LLM texts by calculating the n-gram similarity between the continuation and the original text. Quidwai et al. (2023) proposes calculating cosine similarity to distinguish between the two. While beneficial in that they do not require access to the logits of the source model, black-box statistics methods are challenging due to the substantial overhead of accessing the LLM and long response times.
  • Linguistics Features Statistics. Several recent studies on detecting LLM-generated text have proposed detection methodologies based on linguistics feature statistics. For a survey on features, we refer readers to Beresneva (2016); for a summary of recent research using linguistic features for AI detection, we refer readers to Wu et al. (2023). One limitation of applying linguistic feature statistics is that their availability relies heavily on extensive corpus statistics and various types of LLMs. Our solution of deriving a theoretical bound that guarantees achieving the required level of detection certainty when a small corpus is available can resolve this hurdle.
Possibility of detection & Robustness. Robust detection methods are being developed, and increasingly sophisticated evasion methods are being devised to circumvent these detectors, creating an ongoing contest between detection and evasion. The evasion of detection methods can be through (i) prompt engineering by general users who change the writing style of LLM-generated text or (ii) through paraphrasing attacks by adversaries. Recent research examines the ability of different detection methods against the mentioned evasion methods, including, e.g., Sadasivan et al. 2023; Krishna et al. 2024; Zhang et al. 2024. As a theoretical work in this stream, Chakraborty et al. (2023) provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$. In binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR). For a survey on possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. (2023).