### v3

# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

Abstract
Verifying the provenance of content is crucial to the function of many organizations, e.g., educational institutions, social media platforms, firms, etc. This problem is becoming increasingly challenging as text generated by Large Language Models (LLMs) becomes almost indistinguishable from human-generated content. In addition, many institutions utilize in-house LLMs and want to ensure that external, non-sanctioned LLMs do not produce content within the institution. In this paper, we answer the following question: Given a piece of text, can we identify whether it was produced by LLM $A$ or $B$ (where $B$ can be a human)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design zero-shot statistical tests to distinguish between (i) the text generated by two different sets of LLMs $\mathcal{A}$ (in-house) and $\mathcal{B}$ (non-sanctioned), and also (ii) LLM-generated and human-generated texts. We prove that our tests’ type I and type II errors decrease exponentially as text length increases. In designing our tests, we derive concentration inequalities on the difference between log-perplexity and the average entropy of the string evaluated by $A$. Specifically, for a given string, we demonstrate that if the string is generated by the evaluator model $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in the string length. We also show that if $B$ generates the text, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Lastly, we present two sets of experiments: first, we present experiments using open-source LLMs to support our theoretical results, and then we provide experiments in a black-box setting with adversarial attacks. Practically, our work enables guaranteed (with high probability) finding of the origin of harmful or false LLM-generated text for a text of arbitrary size, which can be useful for combating misinformation as well as compliance with emerging AI regulations.
1 Introduction
LLM text generation tools such as GPT-4 (OpenAI 2023), Llama (Touvron et al. 2023), Gemini (Team et al. 2024) and Mixtral (Jiang et al. 2024) are being widely employed to produce textual content in various domains including news agencies (Newsguard 2024) and academia (Originality.ai 2024). As LLM content generation tools improve, accurately telling whether a text is human-generated or LLM-generated becomes increasingly challenging.
But why do we need to know whether a person or an LLM generated the text? First, LLMs allow people to deliver content they did not produce. In educational settings, for example, this disrupts current accreditation systems. It can also impact the quality of academic peer reviews (Reza et al. 2024). Second, LLMs let people proliferate content. That means an individual can produce significantly more content, including content deliberately tailored to specific audiences, influencing public discussions, e.g., manipulating financial markets (News 2024), political discussions (Times 2023), or consumers’ sentiments (Jakesch et al. (2023)). As stated, not being able to distinguish between LLM and human-generated text yields societal consequences, including the spread of LLM-generated misinformation (Chen and Shu 2024) and LLM-assisted academic cheating Cotton et al. 2024. Therefore, from the policy maker’s perspective, providing methods that reliably distinguish between human and LLM-generated content is crucial. Alongside policymakers, generative AI providers also have incentives to be able to detect AI and human-generated content. That is because retraining on AI-generated content can lead to a phenomenon named "model collapse," first reported by Shumailov et al., 2023. To avoid model collapse, AI providers need to exclude AI-generated content from their training sets, requiring them to identify such content reliably.
In addition to distinguishing human vs. LLM-generated text, henceforth referred to as the detection problem, the ability to differentiate between text generated by a language model vs. another language model, henceforth referred to as the attribution problem, is also critical for the following reasons. First, finding the origin of harmful or false LLM-generated content is essential for legal compliance and mitigation purposes (Hacker et al., 2023; Wang et al., 2023a). Identifying the source of harmful content allows for responsibility to be assigned in case of non-compliance with regulations. Second, the quality of outputs can vary significantly between models, and as a result, employing the most appropriate model for the specific task is essential to achieving optimal results. In line with operating the most suitable language model, educational organizations are building their LLMs to secure educational integrity and credibility. For example, the University of Michigan has developed U-M GPT and UM Maizey as its generative AI tools to ensure academic integrity, guarantee user data protection, and ensure that the shared information does not train the underlying AI models (UMichigan 2023). Students can use only the specialized assistant (here, UM Maizey) to do their assignments, not ChatGPT or any other LLM. This requirement highlights the need for reliable tools to detect the text generated by prohibited LLMs. Again, alongside policymakers, generative AI providers have incentives to be able to detect the text generated by their own language model to gauge the uptake of their systems, which is a commercially important measure of performance.
Figure 1: (Left) LLM-generated Text Detection focuses on detecting whether a text has been generated by an LLM or by a human; (Right) LLM-generated Text Attribution focuses on detecting the LLM source that generated the text among candidate sources.
To evaluate the possibility of detecting LLM-generated text by human linguistics experts, Casal and Kessler, 2023 designed an experiment to investigate whether linguists can distinguish human and ChatGPT-generated text and reported an identification rate of only $38.9\%$. Since humans, even experts, perform poorly in detecting LLM-generated text, researchers are investing significant efforts in designing automated detection methods to identify signals that are difficult for humans to recognize.
One approach to creating a detection method is to train classifiers on labeled training data from LLM-generated and human-generated classes (Bakhtin et al. 2019; Jawahar et al. 2020; Uchendu et al. 2020; Fagni et al. 2021; Sadasivan et al. 2023; Guo et al. 2023; Verma et al. 2023). The limitations of supervised detection approaches make their application practically challenging. First is the requirement for training a separate (from the source-model) classifier, especially considering the large and growing number of LLMs, the wide variety of topics and writing styles, and the possibility of prompting LLMs to write in different styles. Furthermore, the requirement for collecting a dataset of human and AI-generated passages raises concerns, such as privacy, associated with training models on human data. Finally, Liang et al., 2023 notes that because detectors are often evaluated on relatively easy datasets, their performance on out-of-domain samples is often abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using commercial detectors.
An alternative detection method is watermarking (Kirchenbauer et al. 2023), in which an LLM embeds hidden signals into the text during generation, and the signals are identifiable by an algorithm while indistinguishable to humans. Watermarking relies on cooperation from the AI company/owner of the LLM, and current regulations cannot force companies to adopt this technology (Nature, 2024). The described limitations motivate the need for models that do not require training on human data or cooperation from the LLM owner. One such method is zero-shot training, which does not require additional data collection (see, e.g., the work of Mitchell et al., 2023).
The majority of zero-shot detection studies leverage statistical properties of LLM-generated text for identification, including likelihood curvature (Mitchell et al. 2023), log-likelihood (Solaiman et al. 2019), rank-likelihood (Gehrmann et al. 2019), log-likelihood ratio ranking (Su et al. 2023), entropy and Kullback-Leibler (KL) divergence (Lavergne et al. 2008), and perplexity (Vasilatos et al. 2023; Wang et al. 2023b). While most methods focus on a single statistical property, Hans et al. (2024) introduces a normalized measure of perplexity by dividing log-perplexity by average cross-entropy. The rationale is that LLM-generated text is predictable to the LLM (low perplexity), whereas human text is more surprising (higher perplexity). However, high perplexity as a human-authorship signal is unreliable due to prompt dependency, as illustrated by the "Capybara Problem" (Hans et al. 2024). For instance, given the prompt "Can you write about a capybara astrophysicist?", an LLM may generate "The capybara studied dark matter in Andromeda." Without the prompt, the words "capybara" and "astrophysicist" seem unexpectedly high in perplexity, falsely suggesting human authorship. While existing methods, perform well empirically, they remain heuristic-based and lack formal guarantees.
To address this limitation, our study provides a theoretical foundation for distinguishing between LLM-generated and human-written text. In our study, we introduce a framework for determining whether a given string was produced by LLM $A$ or another source $B$ (which may be a human or another LLM). Specifically, for a given string, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. We also show that if $B$ generates the text, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$.
We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single string of text with finite length, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We design tests to distinguish between different LLMs. For this purpose, we assume that we have white-box access to the models in the hypothesis test. In particular, we design composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. We also study the case where we do not have white-box access to all models in the hypothesis set (for example, a human wrote the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
We contribute to the literature on zero-shot statistical tests by developing the first statistical test (with theoretical guarantees) that identifies whether a finite-length text was generated by an LLM or by a human. We show that the type I and type II errors for our statistical tests decrease exponentially as the text length increases.
With the development of specialized LLMs such as UM Maizey, enabling a theoretically guaranteed linking of a finite-length text to its origin among a set of LLMs is becoming necessary. We are the first to provide statistical tests with guarantees for this problem, and we also prove that the type I and type II errors for our statistical test decrease exponentially as the text length increases.
Finally, our theoretical results include establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables. These concentration bounds have been derived for a sequence of independent and identically distributed (iid) random variables by Zhao 2022. However, large language models generate text sequences dependent on the previously generated tokens. Motivated by the problem requirement, we generalize the results in the literature (e.g., Zhao 2022) by proving an exponential decay concentration inequality to bound the tail probability of the difference between the log-likelihood of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables satisfying a martingale structure. Interestingly, as a byproduct of this paper, we address one of the future research directions mentioned by Zhao 2022.
Mudireddy et al. (2024) shows that the log-perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions when the evaluator and the generator model are the same. They note that their result relies on sufficiently large text samples to reach asymptotic behavior. This version of their manuscript, with correct proofs, was made online at the same time/after the initial release of our manuscript. We generalize this work to the case with an arbitrary length text and the case where generator and evaluator models are different by showing the convergence of log-perplexity to cross-entropy between the generator and the evaluator models, except with an exponentially small probability in the string length.
This paper is organized as follows: In Section 2, we present our mathematical framework and define the critical random variables necessary for deriving our theoretical results. Section 3 consists of two subsections. In Subsection 3.1, we provide concentration bounds when the generator and evaluator models are the same, and in Subsection 3.2, we provide concentration bounds when the generator and evaluator models are different. In Section 4, we present our statistical tests and derive the upper bound on their type I and type II errors. In Section 5, we show the results of our experiments on open-source language models. We conclude the paper in Section 6. Appendix A presents the detailed proofs, and Appendix B provides a literature review.
2 Model and background definitions
2.1 Model
Let $M$ be a generative model described by $\textbf{Y}=m(\textbf{X})$, where X denotes user prompt and the output denoted by Y consists of a string of tokens $\textbf{Y}=[Y_{1},Y_{2},\dots,Y_{N},\dots]$. Each token is chosen from a finite vocabulary set, i.e., $Y_{n}\in\mathcal{X}$, and we denote the vocabulary size by $K:=|\mathcal{X}|$.
Practical implementations of LLMs specify the probability distribution iteratively, e.g., Radford 2018. The model first draws a random value for the first token, say $Y_{1}=y_{1}$ by sampling from the distribution $p^{M}(Y_{1}|\textbf{X})$, and then for each token $n\in[2,N]$, the model sequentially determines a distribution for the token given prompt X and all the randomly chosen values $y_{1},y_{2},\dots,y_{n-1}$. So, we define a sequence of probability distributions $p^{M}(\textbf{Y}_{N}|\textbf{X})$ over $\textbf{Y}_{N}\in\mathcal{X}^{N}$ where $\textbf{Y}_{N}=[Y_{1},Y_{2},\dots,Y_{N}]$ is a substring of Y consisting of the first $N$ tokens. The sequence of probability distributions is determined as
$P^{M}(\textbf{Y}_{N}|\textbf{X})=\prod_{n=1}^{N}p^{M}_{n}(\textbf{Y}_{n}),\text{where}\quad p^{M}_{n}(\textbf{Y}_{n})=P^{M}(Y_{n}|Y_{1},Y_{2},\dots,Y_{n-1},\textbf{X}).$ (1)
Note that Equation (1) is an application of the Bayes’ rule and holds for any generative model (regardless of whether tokens $Y_{n}$ are sequentially generated). While equation (1) holds for all generative models, because conditional distributions $p_{n}(y)$ are in general not easily accessible, we apply the rule for sequential models. We follow the literature on white-box detection, and assume that we have complete knowledge of the probability law $p^{M}_{n}(\textbf{Y}_{n})$ for any given sequence $\textbf{Y}_{n}$. See, for example, Mitchell et al. 2023; Gehrmann et al. 2019.
2.2 Background definitions
We aim to design statistical tests to reliably evaluate whether text is generated by model $A$ or by a different model $B$, which can be a different generative model or human.
Cross-entropy. Suppose model $B$ generates a string $\mathbf{Y}$. The cross-entropy of model $B$ and evaluator model $A$ over sub-string $\mathbf{Y}_{N}$ equals
$h_{N}(B,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$ (2)
Intuitively, cross-entropy measures how surprising the token predictions of $A$ are when observed by $B$.
Entropy. Note that if model $A$, is both the generator and the evaluator of the string, then cross-entropy and entropy are equivalent and defined as
$h_{N}(A,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$
Our consideration of white-box detection yields that we have complete knowledge of the probabilities $p^{A}_{n}(y_{n})$, and therefore, we can compute $h_{N}(A,A)$. Furthermore, Gibbs’ inequality (Cover (1999)) states that $h_{N}(B,A)\geq h_{N}(B,B)$.
Perplexity. The perplexity $p^{A}(Y_{N})$ of a (finite length) text string $\textbf{Y}_{N}=[Y_{1},Y_{2},\ldots,Y_{N}]$ with respect to an evaluator model $A$ is defined as the per-token inverse likelihood of the string Y. Formally, perplexity with respect to model $A$ is
$p^{A}(\textbf{Y}_{N})=\left(\prod_{n=1}^{N}p^{A}_{n}(Y_{n})\right)^{-\frac{1}{N}},$
and log-perplexity with respect to model $A$ is
$l_{A}(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\log(p^{A}_{n}(y_{n})).$ (3)
Intuitively, log-perplexity measures how "surprising" a string is to a language model.
2.3 Key random variables
Here, we define the random variables that are critical in deriving our theoretical results.
Let us first define the random variable $Z_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$. If $\mathbb{E}[Z_{n}]<\infty$, then we define a zero-mean random variable $X_{n}=:Z_{n}-\mathbb{E}[Z_{n}]$. If model $A$ generated the string, then we denote the expected value for the random variable $Z_{n}$ by $\mathbb{E}_{p_{n}^{A}}[Z_{n}]$. Lastly, define a random variable $S_{N}:=\sum_{i=1}^{N}X_{i}$.
3 Concentration bounds
In this section, we present our results in two parts. In Section 3.1, we provide concentration bounds to show that if the string is generated by model $A$, $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average entropy of the string under $A$ with a high probability. In Section 3.2, we provide concentration bounds to show that if the string is generated by another model $B$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$ converges to the average cross-entropy of the string under $B$ and $A$ with a high probability. These concentration bounds are the backbones of the statistical tests that we design in section 4.
3.1 Same generative and evaluator models
Consider a string $\mathbf{Y}$ generated by model $A$ and we evaluate the text using the same model $A$. First recall that given a string $\mathbf{Y}_{N}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$. Then,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n}).$
Also, given the string $\mathbf{Y}_{N}$, since we have complete knowledge on the probabilities $p^{A}_{n}(y_{n})$, we have complete knowledge on the entropy $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})=\mathbb{E}_{p_{n}^{A}}[Z_{n}]$. Then, we can find the following upper bound for $\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Remark 1.
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})\leq\log|\mathcal{X}|=\log(K).$
Proof.
By concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, the value of its maximizer is $p^{*A}_{n}(y_{n})=\frac{1}{|\chi|},\quad\forall y_{n}\in\chi$. Thus,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}]=-\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log p_{n}^{A}(y_{n})$$\leq$ $\sum_{y_{n}\in\chi}p_{n}^{*A}(y_{n})\log p_{n}^{*A}(y_{n})$
$=$ $|\chi|.\frac{1}{|\chi|}.\log|\chi|=\log(K).$
With $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p_{n}^{A}}[Z_{n}]$.
Lemma 1.
The random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale.
Proof.
$Z_{n}$ and as a result $\mathbb{E}_{p_{n}^{A}}(y_{n})$ are positive random variables, and we showed that $\mathbb{E}_{p_{n}^{A}}[Z_{n}]<\log(K)$. Therefore, we have $\mathbb{E}\big{[}|X_{n}|\big{]}\leq\log(K)$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}|S_{N}]=S_{N}$.
(i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}-S_{N}|\big{]}=\mathbb{E}\big{[}|X_{N+1}|\big{]}\leq\log(K)<\infty$.
(ii) is satisfied because the martingale increments $X_{n}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Finally, we apply concentration bounds for martingales to provide finite sample guarantee for the convergence of the random variable $S_{N}/N$ to zero. A challenge in applying the common concentration bounds for martingales is that martingale increments are not necessarily bounded. We overcome this issue by showing that the martingale differences, while not bounded, admit a light tail. In particular, we show that the martingale differences are sub-exponential.
Definition 1.
(sub-exponential norm). The sub-exponential norm of $X\in\mathbb{R}$ is
$\|X\|_{\psi_{1}}=\inf\bigg{\{}t>0:\mathbb{E}[e^{\frac{|X|}{t}}]\leq 2\bigg{\}}.$
If $\|X\|_{\psi_{1}}$ is finite, we say that $X$ is sub-exponential.
Lemma 2.
The sub-exponential norm for random variable $X_{n}$ equals $2\log(K)$.
Proof.
See Appendix A.1. ∎
As the last step, we apply the concentration bounds for martingales with sub-exponential increments to obtain the following concentration bound.
Theorem 1.
There exists a constant $c_{1}>0$ independent of the evaluator model $A$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}$
Proof.
See Section A.2. ∎
Interpretation. Theorem 1 states that if a given string is generated by a model same as the evaluator (here, model $A$), then the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length.
3.2 Different generative and evaluator models
In this section, we consider a string $\bf{Y}$ generated by model $B$ and want to evaluate our statistical test based on model $A$. To design the statistical test, first recall that given a string $\bf{Y}$ and an evaluator model $A$, we define the random variable $Z_{n}=-\log\big{(}p^{A}_{n}(y_{n})\big{)}$. Note that
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))=H(p^{B}_{n},p^{A}_{n}),$ (4)
where $H(p^{B}_{n},p^{A}_{n})$ is the cross entropy between the two distributions $p^{B}_{n}(.)$ and $p^{A}_{n}(.)$.
Note that, unlike the case analyzed in section 3.1, here, $\mathbb{E}_{p_{n}^{B}}[Z_{n}]$ is not necessarily finite.
For $\mathbb{E}_{p^{B}_{n}}[Z_{n}]$ to be infinite, as we can infer from equation 4, we must have that $p^{A}_{n}(y_{n})=0$ and $p^{B}_{n}(y_{n})>0$ for some $y_{n}\in\chi$. In this case, if the string includes such $y_{n}$, then we realize that the string is not generated by model $A$ with the probability of 1. This is a trivial case.
Yet, if the string does not include any such $y_{n}$, then we can update the probability distributions as
$\tilde{p}^{B}_{n}=\frac{p^{B}_{n}(y_{n})}{\sum_{y_{k}:p^{A}_{n}(y_{k})>0}p^{B}_{n}(y_{k})}.$
Without loss of generality, we can exclude the trivial case and assume that if $p^{B}_{n}(y_{n})>0$, then $p^{A}_{n}(y_{n})>0$, which yields that $\mathbb{E}_{p^{B}_{n}}[Z_{n}]=H(p^{B}_{n},p^{A}_{n})$ is finite. With $\mathbb{E}_{p_{n}^{B}}[Z_{n}]<\infty$, we define the zero-mean random variable $X_{n}=Z_{n}-\mathbb{E}_{p^{B}_{n}}[Z_{n}]$. For tractability, we make a parametric assumption on the probability laws $p^{A}_{k}(.)$ and $p^{B}_{k}(.)$.
Assumption 1.
For our two model analysis, we assume that there exists $\epsilon>0$ such that $p^{A}_{n}(y_{k}),p^{B}_{n}(y_{k})\notin(0,\epsilon)$.
Assumption 1 implies that models $A$ and $B$ either do not associate any probability to a token $y\in\mathcal{X}$, or they assign a probability of at least $\epsilon$. Our theoretical results depend only on $\log(\epsilon)$. Hence, our theoretical bounds primarily rely on a constant shift in the logarithmic scale.
It is noteworthy that Assumption 1 is not restrictive and is aligned with practice, as computers only allow for a limited range of representable numbers due to finite precision in floating-point arithmetic. Very small probabilities are either rounded to zero or set to a minimum threshold to maintain numerical stability in computations (Goldberg, 1991).
The first outcome of assumption 1 is that
$\mathbb{E}_{p^{B}_{n}}\big{[}|Z_{n}|\big{]}=\mathbb{E}_{p^{B}_{n}}\big{[}Z_{n}\big{]}=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon)$ (5)
Hence, with the same argument as in section 3.1, the random variable $S_{N}=\sum_{i=1}^{N}X_{i}$ forms a martingale. To apply the martingale concentration bounds, similar to the previous section, we first find the sub-exponential norm for the random variable $X_{n}$.
Lemma 3.
Under Assumption 1, the sub-exponential norm for the random variable $X_{n}$ equals $-4\log(\epsilon)$.
Proof.
See Section A.3. ∎
Theorem 2.
There exists a constant $c_{3}>0$ independent of models $A$ and $B$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}$
Proof.
See Section A.4. ∎
Interpretation. Theorem 2 states that if model $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under model $A$ converges to the average cross-entropy of the string under $B$ and $A$.
4 Statistical test
Now we design our statistical tests using the results in Theorems 1 and 2 and then evaluate type I (false positive) and type II (false negative) errors. In particular, we consider a finite-length text with length $N$ generated by a model $M$. We first design tests for detection between different LLMs. In that, we have white-box access to all models in the hypothesis test. In Section 4.1, we design a simple statistical test that determines whether a text is generated by a model $A$ or another model $B$. Then, in Section 4.2, we extend our results to composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. Finally, in Section 4.3, we study the case where we don’t have white-box access to all models in the hypothesis set (for example, a human produced the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
4.1 Simple statistical test for detection between two LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with length $N$, we design a statistical test to detect whether model $A$ or model $B$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by $B$, and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by $A$. We first calculate the random variables $Z^{A}_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$ and $Z^{B}_{n}=:-\log\big{(}p^{B}_{n}(Y_{n})\big{)}$, and then we calculate the sums $\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}$, and $\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the model $A$ and incorrectly concludes that it is generated by $B$.
To quantify our model’s type I and type II errors, we need to make the following (mild) assumption.
Assumption 2.
(minimum difference). We assume that if the generative and evaluator models are different, for an arbitrarily small positive $\epsilon_{1}>0$, we have
$\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\geq\epsilon_{1}.$
Assumption 2 ensures that the two models satisfy a minimum distance in terms of their KL divergence over the generated text. Note that KL divergence, by definition, is a non-negative value that demonstrates the distance between the two distributions over the next word for the two models. Our results show that the type I and type II errors of our statistical test are approximately $\exp\big{(}O(-N\epsilon_{1})\big{)}$, which indicates that even for small values of $\epsilon_{1}$ that can converge to zero with the length of text (for example, $\epsilon_{1}=O(N^{-1/2})$), our statistical test provides exponentially small type I and type II errors in the length of the text. Hence, our theoretical bounds only require that the two models do not impose the same probability distribution on the string.
Proposition 1.
If Assumptions 1 and 2 hold, then the type I and type II errors for our statistical test are upper bounded by
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.5. ∎
Interpretation. Proposition 1 demonstrates that the type I and type II errors of our simple test decrease exponentially in the text length.
4.2 Statitical test for detection among multiple LLMs
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, we design a statistical test to detect whether the text is generated by one of the models $\mathcal{A}=\{A_{1},\ldots,A_{p}\}$ or one of models =$\mathcal{B}=\{B_{1},\ldots,B_{q}\}$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by one of the models in $\mathcal{B}$, and the alternative hypothesis $\mathbf{H}_{1}$ is that it is generated by one of the models in $\mathcal{A}$. We first calculate the random variables $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, and sum $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, for all models $M\in\mathcal{A}\cup\mathcal{B}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ if for some $A_{i}\in\mathcal{A}$, we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n};\qquad\forall B_{j}\in\mathcal{B}.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by one of the models in $\mathcal{A}$ when it is written by one of the models in $\mathcal{B}$, and Type II error happens when the test fails to identify that text is generated by one of the models in $\mathcal{A}$ and incorrectly concludes that it is generated by one of the models in $\mathcal{B}$. Similar to our test for the two model version, Assumption 2 must hold for us to quantify our model’s type I and type II errors.
Proposition 2.
If Assumptions 1 and 2 hold, the type I error for our statistical test is upper bounded by
$2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
and the type II error for our statistical test is upper bounded by
$2|\mathcal{B}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorems 1 and 2.
Proof.
See Section A.6. ∎
Interpretation. Proposition 2 demonstrates that the type I and type II errors of our composite test decrease exponentially in the text length.
4.3 Statistical test for detection between an LLM and human
We need to make the following assumption to design our test.
Assumption 3.
(minimum tangible difference). We assume that if the generative and evaluator models are different, then
$\frac{1}{N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}\textbf{Y}_{n-2}\bigg{]}\geq 4\log^{2}\big{(}K\big{)}.$
Assumption 3 ensures that the two models satisfy a minimum distance in terms of their expected KL divergence. Clearly, if models are the same (or very similar), then $A$ and $B$ impose the same (or almost the same) probability distributions $p_{N}^{A}(y_{n})$ over $\mathbf{Y}_{N}$, and hence KL divergence becomes zero. This makes differentiation impossible. Assumption 3 rules out cases where the two models impose very similar distributions over the text under evaluation.
Lemma 4.
Under assumption 3, for any positive constant $c_{4}\leq\log^{2}(K)/2$, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Proof.
See Section A.7. ∎
Remark 2.
We note that the only application of Assumption 3 for establishing our results is that this assumption ensures that the statement $c_{4}\leq\log^{2}(K)/2$ in Lemma 4 holds. Our results remain true ( even without Assumption 3) if there exists $c_{4}>0$ such that if the generative and evaluator models are different, then
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Statistical test. Given a string $\mathbf{Y}_{N}$ with size $N$, for arbitrary constants $t<c_{4}\leq\log^{2}(K)/2$, we design a statistical test to detect whether the evaluator model $A$ generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is not generated by the evaluator model $A$ (e.g., it is generated by another model $B$), and the alternative hypothesis $\mathbf{H}_{1}$ is that $\mathbf{Y}_{N}$ is generated by the evaluator model $A$. We first calculate the random variable $Z_{n}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and then we calculate the sum $\frac{1}{N}\sum_{n=1}^{N}Z_{n}$. Our test rejects the null hypothesis $\mathbf{H}_{0}$ in favor of the alternative $\mathbf{H}_{1}$ if
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Otherwise, our test accepts the null hypothesis.
Type I and type II errors. Type I error occurs when the test incorrectly concludes that the text is generated by the evaluator model $A$ when it is written by $B$, and Type II error happens when the test fails to identify that text is generated by the evaluator model $A$ and incorrectly concludes that it is not written by $A$.
Proposition 3.
For any $t\geq 0$, the type II error for our statitical test is upper bounded by
$2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
with $c_{1}$ as introduced in Theorem 1.
Additionally, if Assumptions 1 and 3 hold, then for any positive $t<c_{4}$, type I error for our statistical test is upper bounded by
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.,$
with $c_{3}$ as introduced in Theorem 2 and $c_{4}$ as introduced in Assumption 3.
Proof.
See Section A.8. ∎
Interpretation. Proposition 3 demonstrates that the type I and type II errors of our test for detecting LLM $A$ vs. not LLM $A$ (that includes human) decrease exponentially in the text length.
5 Experiments
In this section, we present two main sets of experiments: (i) experiments with white-box access to the conditional probability distributions for token generation to examine our theoretical results, and (ii) experiments that relax the white-box access assumption. In the black-box setting, we also examine robustness to decoding strategies and adversarial attacks.
Figure 2: Overview of the experiments
Lastly, we provide experiments to demonstrate that the log-perplexity under the evaluator model and the average cross-entropy between the generative and the evaluator model converge.
5.1 White-box detection
Datasets. We use the datasets used in Mitchell et al. (2023) for our experiments. That includes the news articles from the XSum dataset (Narayan et al. 2018), prompted stories from Reddit’s WritingPrompts dataset (Fan et al. 2018), and Wikipedia paragraphs from the SQuAD contexts (Rajpurkar 2016) that contain human-written questions and machine-generated answers from passages from Wikipedia articles. Each experiment uses 150 to 500 examples for evaluation. We prompt the first 30 tokens of the actual text to the LLM to generate machine-generated text and obtain the generated text. For QA datasets, we prompt the question. In all experiments, a positive signal is that the focal LLM generates text, and all experiments use an equal number of positive and negative examples. The language models we examine for whitebox detection are LLAMA 3 (8B parameters), GPT-NEOX Erebus (20B parameters), and QWEN (32B parameters). We use AUROC and TPR @ FPR=1%, which follows the recent research that uses this evaluation paradigm (Krishna et al. 2024; Hans et al. 2024; Soto et al. 2024) as our performance metrics.
AUROC. The Area Under the Receiver Operating Characteristic (ROC) Curve is a standard widely adopted detection performance measure, e.g., Verma et al. 2023; Mitchell et al. 2023, and we use this metric consistent with the literature. The ROC curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at all possible classification thresholds, showing the tradeoff between correctly detecting LLM-generated text and mistakenly flagging human-written text.
TPR @ FPR=1%. Results of experiments in the literature, e.g., in Dugan et al. 2024, suggest that the false positive rates of detectors can be dangerously high, which makes them unfit for regulatory actions. To show the validity of our theoretical results, we examine the accuracy captured by the TPR while maintaining the low FPR.
Benchmarks. We compare the performance of our method with various existing zero-shot detection methods that similarly leverage the predicted token-wise conditional distributions of the source model. The benchmarks we use apply detection tests based on token log-probabilities, token (log-)ranks, likelihood curvature, or (a measure of) predictive entropy (Gehrmann et al. 2019; Solaiman et al. 2019; Ippolito et al. 2019; Mitchell et al. 2023; Hans et al. 2024). Detection based on token log-probabilities uses the source model’s average token-wise log-probability. Detection based on token (log-)rank uses the average observed (log-)rank of the tokens in the passage based on the model’s conditional distributions. The entropy-based detection benchmark inspired by Gehrmann et al. 2019 and applied by Mitchell et al. 2023 hypothesizes that machine-generated texts are more "in-distribution" for the model, leading to lower entropy predictive distributions. Detection based on Mitchell et al. 2023’s framework states that a machine-generated text is sampled mostly at the mode of its text distribution, while human texts might lay anywhere on the text distribution. Lastly, Hans et al. (2024) applies a normalized measure of perplexity by dividing log-perplexity by average cross-entropy.
5.2 White-box detection results
AUROC. In these experiments, model samples are generated by sampling from the raw conditional distribution with temperature 1.
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.99* 0.84 0.99* 0.91 0.75 0.60 1.00 0.95 1.00*
Rank 0.75 0.69 0.69 0.71 0.71 0.55 0.82 0.80 0.81
LogRank 0.99* 0.87 0.99* 0.93* 0.81 0.62 1.00 0.97* 1.00*
Entropy 0.39 0.70 0.40 0.37 0.66 0.70* 0.04 0.36 0.02
DetectGPT 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Binoculars 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98* 1.00 1.00
Diff (zero-shot) 0.00 0.04 0.01 0.06 0.18 0.27 -0.02 0.03 0.00
Roberta (base) 0.98* 0.95 0.92* 0.97* 0.92 0.69* 0.97* 0.95* 0.74*
Roberta (large) 0.98* 0.98* 0.92* 0.95 0.93* 0.68 0.96 0.93 0.65
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98 1.00 1.00
Diff (supervised) 0.01 0.01 0.08 0.02 0.06 0.28 0.01 0.05 0.26
Table 1: AUROC for detecting samples from the given models on the datasets using our and other baseline methods. The best AUROC values are in bold, and the second-best values are marked with an asterisk (*). The rows Diff(zero-shot) and Diff(supervised) show our AUROC improvement over the strongest zero-shot and supervised baseline methods.
Our method achieves its highest improvement in detecting LLM-generated text for SQuAD, increasing AUROC from 0.04 to 0.27. It also improves detection for XSum (0.015 AUROC) and WritingPrompts (0.025 AUROC), though with a smaller margin. An explanation for the weaker performance of existing detectors on SQuAD is its higher entropy variance (0.128) compared to XSum (0.045) and WritingPrompts (0.074). Since our method relies on the difference between log-perplexity and empirical entropy, it performs more effectively when entropy variance is high while aligning with likelihood-based methods when entropy is less variable. Finally, our method significantly outperforms existing likelihood-based approaches in detecting Qwen-generated text for SQuAD.
TPR @ FPR=1%. In Table 2, we show the accuracy captured by TPR, while the FPR is maintained low (1%).
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.89 0.02 0.99* 0.01 0.05 0.01 1.00 0.46 1.00
Rank 0.22 0.06 0.20 0.07 0.18 0.01 0.49 0.39 0.43
LogRank 0.99 0.05 0.99* 0.02 0.06 0.01 1.00 0.54 1.00
Entropy 0.02 0.09 0.01 0.01 0.04 0.00 0.00 0.00 0.00
Binoculars 0.99 0.89* 0.99* 0.96 0.86* 0.26* 0.97* 0.99 0.99
Ours 0.97* 1.00 1.00 0.96 0.92 0.72 0.90 0.99 1.00
Diff (zero-shot) -0.02 0.11 0.01 0.00 0.06 0.46 -0.10 0.00 0.00
Table 2: TPR at FPR=0.01. The best TPR values are in bold, and the second-best values are marked with an asterisk (*). The row Diff(zero-shot) shows our TPR improvement over the strongest zero-shot baseline methods.
The experiment’s results show the notable superiority of our method in fact-based datasets (e.g., SQuAD), where the human-written and LLM-generated texts tend to be similar, resulting in lower performance of conventional detection methods. Specifically, we observe the maximum gap between ours and the second-best methods in the SQuAD dataset with QWEN 32B as the evaluator. Even when the FPR is held low, the human and LLM texts are more distinguishable in datasets with less fact-based and more creative content. Also, smaller language models (e.g., LLAMA 2 8B) are easier to detect.
5.3 Black-box detection
Datasets. We use the datasets in Dugan et al. 2024 that consist of data in 8 domains of abstracts, books, news, poetry, recipes, Reddit, reviews, and Wikipedia. The non-adversarial dataset consists of 14,971 human-written from publicly-available pre-2022 datasets and 509,014 LLM-generated documents for a total of 6,287,820 texts when including adversarial attacks.
Benchmarks. We compare the performance of our method with the detectors in Dugan et al. 2024 that include neural detectors, zero-shot detectors, and commercial detectors. The neural detectors are RoBERTa-B (GPT-2), RoBERTa-L (GPT-2), RoBERTa-B (ChatGPT), and RADAR. The zero-shot detectors are GLTR, Fast DetectGPT, Binoculars, and LLMDet. The commercial detectors are GPTZero, Originality, Winston, and ZeroGPT. In this section, we do not have white-box access to any of the detectors, and we use tiiuae/falcon-rw-1B as our scoring model.
Repetition penalty. This penalty down-weighs the probability of tokens that previously appeared in the context window by some multiplicative factor $\theta$, resulting in less repetitive output. Following Keskar et al. 2019, we use $\theta=1.2$ for our experiments.
Adversarial attacks. When selecting adversarial attacks, consistent with Dugan et al. 2024, we assume that the adversary uses predefined transformations or heuristics that degrade typical detection cues, different from the gradient-based methods. Accordingly, we include the following 8 black-box, query-free attacks:
  1. 1.
    Alternative Spelling (AS): uses British spelling instead of the American spelling.
  2. 2.
    Article Deletion (AD): deletes "the", "a", "an."
  3. 3.
[…]
  4. 4.
    Number Swap (NS): randomly shuffles number digits.
  5. 5.
    Misspelling (MS): inserts common misspellings.
  6. 6.
    Synonym Swap (SS): swaps tokens with highly similar BERT (Devlin et al. (2019)) candidate tokens.
  7. 7.
    Upper Lower Swap (ULS): swaps the case of words.
  8. 8.
    Whitespace Addition (WSA): adds spaces between characters.
Note that two other adversarial attacks, i.e., Homoglyph and Zero-width space, are possible. Homoglyph swaps characters in a text with visually similar yet technically different characters, often from other alphabets or Unicode blocks. Zero-width space inserts a Unicode character (often U+200B) that does not occupy visible space. Dugan et al. (2024)’s dataset for these attacks was incomplete, not allowing us to provide results for these attacks. However, these two attacks are easy to remove with the following pre-processing:
  • Removing or normalizing invisible characters like zero-width spaces
  • Mapping suspicious non-ASCII characters to standard ASCII or a canonical form
5.4 Black-box detection results
Accuracy at FPR=5%. Consistent with Dugan et al. (2024), we evaluate accuracy by TPR while fixing FPR at 5%. In the experiments, the open-source chat models include Llama-c, Mistral-c, and MPT-c. The open-source non-chat models include Mistral, MPT, and GPT2. The closed-sourced chat models include c-GPT, GPT4, and Cohere, and the closed-sourced non-chat models include Cohere and GPT3. For the repetition penalty, we set the multiplicative factor as $\theta=1.2$.
Open-Source Closed-Source
Chat Models Non-Chat Models Chat Models Non-Chat Models
Rep. Penalty? × × × ×
GPTZero 98.4 82.5 9.4 4.8 88.5 53.4
Originality 97.7 72.5 89.0 51.2 89.0 85.4
Winston 96.6 78.3 29.5 11.3 93.7 68.1
ZeroGPT(*) 90.5 54.9 16.0 0.3 65.8 72.7
R-B GPT2 77.9 26.2 60.5 35.4 41.7 52.5
R-L GPT2 71.4 19.5 67.2 53.4 34.7 48.6
R-B CGPT 75.0 39.3 14.9 1.7 38.1 39.0
RADAR 85.6 66.4 48.3 31.8 75.3 67.7
GLTR 83.9 38.3 44.5 0.5 54.3 63.7
F-DetectGPT 96.2 40.5 79.7 0.6 74.1 86.3
LLMDet 47.5 16.5 38.4 3.7 18.5 32.9
Binoculars 99.7 60.6 72.4 0.6 92.1 95.0
Ours 97.6 60.7 51.7 71.2 68.8 86.5
Table 3: Accuracy Score at FPR=5% for all detectors across model groups and sampling strategies. Asterisks (*) indicate that the detector was unable to achieve the target FPR. For accuracy of 67%-100%, we highlight green, 33%-66% we highlight yellow, and for 0%-32% we highlight red.
For open-source chat models without a repetition penalty, our method achieves an accuracy of 97.6%, indicating that when text maintains its natural stylistic and structural cues, our detector can accurately flag LLM-generated content. For open-source non-chat models, the addition of a repetition penalty improves performance, with accuracy rising from 51.7% without the penalty to 71.2% with the penalty. In the closed-source setting, our detector again performs robustly—68.8% for chat models and a high 86.5% for non-chat models, demonstrating our method’s capacity to generalize across different text styles. Overall, these findings indicate that our detector is particularly strong when the underlying text generation preserves the natural language features that our method is designed to leverage, and that applying a repetition penalty can further enhance accuracy in contexts where repetitive patterns might otherwise obscure these signals.
Adversarial attacks
None AS AD IP NS MS SYN ULS WSA
RoB-B GPT2 59.1 55.6 37.1 56.9 55.9 43.8 71.5 18.8 45.2
RoB-L GPT2 56.7 52.4 33.2 55.1 51.7 39.5 79.4 19.3 40.1
RoB-B CGPT 44.8 43.3 38.0 5.2 44.3 42.1 39.6 31.7 0.1
RADAR 70.9 70.8 67.9 73.7 71.0 69.5 67.5 70.4 66.1
GPTZero 66.5 64.9 61.0 66.2 65.8 65.1 61.0 56.5 66.2
ZeroGPT 65.5 65.4 59.7 64.9 64.7 64.7 18.8 54.5 64.2
Originality 85.0 83.6 71.4 85.1 86.0 78.6 96.5 75.8 84.9
Winston 71.0 68.9 66.9 69.8 69.0 67.5 63.6 56.8 46.8
GLTR 62.6 61.2 52.1 61.4 59.9 59.8 31.2 48.1 45.8
F-DGPT 73.6 71.6 64.7 72.0 68.2 70.7 34.0 60.4 64.4
LLMDet 35.0 33.9 27.4 27.2 33.8 32.7 27.3 23.4 4.4
Binoculars 79.6 78.2 74.3 71.7 77.1 78.0 43.5 73.8 70.1
Ours 84.3 79.9 73.1 80.7 73.3 78.1 48.0 72.9 57.1
Table 4: Accuracy Score at FPR=5% for all detectors across different adversarial attacks. Abbrevations are: AS: Alternative Spelling, AD: Article Deletion, IP: Insert Paragraphs, NS: Number Swap, MS: Misspelling, SYN: Synonym Swap, ULS: Upper Lower Swap, WSA: Whitespace Addition. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
Our detector maintains relatively high accuracy across multiple adversarial scenarios. In the baseline “None” condition, it records an accuracy of 84.3%. Under attacks such as Alternative Spelling (79.9%), Article Deletion (73.1%), Insert Paragraphs (80.7%), Number Swap (73.3%), Misspelling (78.1%), and Upper–Lower Swap (72.9%), the accuracy remains in the 70–80% range. However, similar to other zero-shot detectors, our performance is reduced under Synonym Swap (48.0%) and Whitespace Addition (57.1%). This detection performance drop is predictable because zero-shot detection methods often rely on fixed token distributions. Synonym substitutions replace surface-level tokens with semantically equivalent but lexically distinct items, thereby shifting the token distribution away from the detector models’ reference distribution. Whitespace manipulations alter standard sub-word segmentation by inserting or removing token boundaries, creating a discrepancy between the text’s actual token frequency and what the detector expects.
5.5 Convergence results
We conduct our numerical analysis for convergence on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
Same generative and evaluator model. In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (3(a)-3(d)). We consistently observe that the numerical results confirm Theorem 1 that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
(a) GPT-2 small
(b) GPT-2 medium
(c) GPT-2 large
(d) GPT-2 XL
Figure 3: Generated and evaluated by the same model
Different generative and evaluator models. To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (4(a)-4(c)). Results in these figures confirm Theorem 2. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XL
Figure 4: Generated by GPT-2 small and evaluated by different models
6 Conclusion
In this study, we establish first zero-shot statistical tests with theoretical guarantees for text with finite length to distinguish between (i) LLM-generated and human-generated texts in Proposition 3, and (ii) the text generated by two different LLMs $A$ and $B$ in Propositions 1 and 2. We prove that the type I and type II errors for our tests decrease exponentially in the text length. As a critical step in designing our tests, we derive concentration bounds in the difference between log-perplexity and the average entropy of the string under $A$. Specifically, for a given string, in Theorem 1, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. Furthermore, in Theorem 2, we show that if $B$ (which can be either another model or human) generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Our theoretical results rely on establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables and the negative entropy for non-independent random variables on a finite alphabet. Results in the literature (e.g., Zhao (2022)) derive concentration bounds for iid random variables, and one of our theoretical contributions is to extend the results to non-independent random variables by introducing random variables that form a martingale. We hope that our work inspires more research on zero-shot LLM-text detection with provable guarantees.
Appendix A Appendix
A.1 Proof of Lemma 2
Proof.
For the random variable to be sub-exponential, by definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\frac{1}{p^{A}_{n}(y_{n})^{1/t}}=e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq 2.$ (6)
where the first inequality follows from the triangle inequality and that $Z_{n}$ and therefore $\mathbb{E}[Z_{n}]$ are positive. Note that $\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}$ is concave in $p^{A}_{n}(y_{n})$. Hence, it attains its maximum when we have $p^{A}_{n}(y_{i})=p^{A}_{n}(y_{j})\quad\forall y_{i},y_{j}\in\mathcal{X}$, which yields
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq K(\frac{1}{K})^{\frac{t-1}{t}}=K^{\frac{1}{t}}.$ (7)
To analyze $e^{\frac{\mathbb{E}[Z_{n}]}{t}}$, we first want to show that
$\mathbb{E}[Z_{n}]$ $=$ $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ (8)
$\leq$ $\log|\mathcal{X}|=\log(K).$
To show that, note that by concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, we can maximize $p_{n}(y_{n})\log p_{n}(y_{n})$ by equalizing all probabilities $p^{*}(y_{n})=\frac{1}{|\mathcal{X}|};\forall n$. So, we have
$-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$$\leq$ $-\sum_{y_{n}\in\mathcal{X}}p^{*}(y_{n})\log p^{*}(y_{n})$
$=$ $-|\mathcal{X}|\Big{[}\frac{1}{|\mathcal{X}|}\log\frac{1}{|\mathcal{X}|}\Big{]}=\log|\mathcal{X}|.$
Combining equations (6)-(8) yields that for $X_{n}$ to be sub-exponential with norm $t$, it must satisfy
$K^{1/t}K^{1/t}\leq 2$
Hence, $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=2\log(K)$. ∎
A.2 Proof of Theorem 1
Proof.
We first present an equivalent (up to a constant factor) definition of a sub-exponential random variable from Vershynin (2018).
Definition 2.
(sub-exponential random variable). Centered random variable $X\in SE(\nu^{2},\alpha)$ with parameters $\nu,\alpha>0$ is sub-exponential if
$\mathbb{E}[e^{\lambda X}]\leq e^{\frac{\lambda^{2}\nu^{2}}{2}},\quad\forall\lambda:|\lambda|<\frac{1}{\alpha}.$
Next, we present a lemma from Vershynin (2018) that demonstrates that the two definitions are equivalent up to a constant factor.
Lemma 5.
(SE properties ,Vershynin, 2018). Let $X$ be a random variable with $\mathbb{E}[X]=0$. Then, there exists a constant $c$ and constants $K_{4}$ and $K_{5}$ such that $K_{4}\leq cK_{5}$ and $K_{5}\leq cK_{4}$ and the following two properties are equivalent.
  • There exists a constant $K_{4}$, such that the MGF of $|X|$ is bounded, specifically
    $\mathbb{E}[e^{|X|/K_{4}}]\leq 2.$
  • There exists a constant $K_{5}$, such that the MGF of $X$ satisfies
    $\mathbb{E}[e^{\lambda X}]\leq e^{\frac{K_{5}^{2}\lambda^{2}}{2}}\quad\forall\lambda\quad s.t.\quad|\lambda|\leq\frac{1}{K_{5}}$
Next, note that by Lemma 2 and Lemma 5, there exists $c_{1}>0$ s.t. for $\nu=c_{1}\log K$, and for any $n\in\mathbb{N}$,
$\mathbb{E}[e^{\lambda X_{n}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (9)
Let $\alpha=\nu$. By Definition 2, we have $X_{n}\in SE(\nu^{2},\alpha)$.
Our next step is to show that $\sum_{n=1}^{N}X_{n}$ is SE with parameters $(\nu\sqrt{N},\alpha_{*})$. To realize that, observe that
$\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n}X_{k}\right)}\right]=\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n-1}X_{k}\right)}\mathbb{E}\left[e^{\lambda X_{n}}\mid\mathbf{Y}_{n-1}\right]\right]\leq\mathbb{E}\left[e^{\lambda\sum_{k=1}^{n-1}X_{k}}\right]e^{\frac{\lambda^{2}\nu^{2}}{2}}\leq e^{\frac{\lambda^{2}N\nu^{2}}{2}},$
where the first equation follows from the iterated law of expectation, and the first inequality follows from Equation (9).
Finally, from Theorem 5.2 in Arinaldo, 2019, if $S\in SE(\nu^{2},\alpha)$ is a sub-exponential random variable, then
$\mathbb{P}(|S-\mathbb{E}[S]|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{\nu^{2}},\frac{t_{1}}{\alpha}\bigg{)}\bigg{)}.$ (10)
Substituting $\nu=\alpha\sqrt{N}$ for the zero-mean random variable $\sum_{n=1}^{N}X_{n}$, and $\alpha=c_{1}\log K$, we obtain
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(c_{1}\log K)^{2}},\frac{t_{1}}{c_{1}\log K}\bigg{)}\bigg{)}.$
Setting $t_{1}=tN$, we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{i=1}^{N}X_{i}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.3 Proof of Lemma 3
Proof.
For the random variable to be sub-exponential, by definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{p^{A}_{n}(y_{n})^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{\epsilon^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}}\leq 2.$
Thus, for $X_{n}$ to be sub-exponential with norm $t$, it is sufficient to satisfy
$\frac{\mathbb{E}[Z_{n}]-\log(\epsilon)}{t}\leq\log(2)\leq 1/2.$ (11)
Recall equation 5 that states
$\mathbb{E}_{p^{B}_{n}}[Z_{n}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon).$
Substituting the result of (5) in (11) yields that $X_{n}$ has sub-exponential norm $\|X\|_{\psi_{1}}=-4\log(\epsilon)$. ∎
A.4 Proof of Theorem 2
Proof.
Following the same steps as in proof of Theorem 1, we conclude that $\sum_{n=1}^{N}X_{n}$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t_{1}}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Finally, setting $t_{1}=tN$, we obtain
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
A.5 Proof of Proposition 1
Proof.
Type I error occurs if the model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
By triangle inequality, this yields
$h_{N}(B,A)-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq h_{N}(B,B)-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B,A)-h_{N}(B,B)\geq\epsilon_{1}$. Hence, the type I error occurs only if
$\epsilon_{1}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Equivalently, the type I error occurs only if
$\epsilon_{1}\leq\bigg{|}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\bigg{|}+\bigg{|}-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}\bigg{|}.$
So, the type I error only occurs if at least $|-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 2, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
The type II error occurs if the model $A$ generates the text $\bf{Y}_{N}$, but we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}.$
By symmetry, the type II error adheres to the same upper bound as the type I error, yielding
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.6 Proof of Proposition 2
Proof.
Type I error occurs if a model $B_{j}\in\mathcal{B}$ generates the text string $\mathbf{Y}_{N}$, but for one model $A_{i}\in\mathcal{A}$ we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
This yields for some $A_{i}\in\mathcal{A}$, we have
$h_{N}(B_{j},A_{i})-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq h_{N}(B_{j},B_{j})-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B_{j},A_{i})-h_{N}(B_{j},B_{j})\geq\epsilon_{1}$. Hence, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
Equivalently, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}+\bigg{|}-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}.$
So, the type I error only occurs if, for at least one of the models $A_{i}\in\mathcal{A}$, we have either $|-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+|\mathcal{A}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 2, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}+2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if a model $A_{i}\in\mathcal{A}$ generates the text $\bf{Y}_{N}$, but for a model $B_{j}\in\mathcal{B}$ we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
This yields that for some $B_{j}\in\mathcal{B}$, we have
$h_{N}(A_{i},B_{j})-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq h_{N}(A_{i},A_{i})-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Note that by Assumption 2, we have $h_{N}(A_{i},B_{j})-h_{N}(A_{i},A_{i})\geq\epsilon_{1}$. Hence, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Equivalently, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}+\bigg{|}-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}.$
So, the type II error only occurs if at least for one of the models $B_{j}\in\mathcal{B}$, we have either $|-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type II error as
$|\mathcal{B}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 2 we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Also, from Theorem 1 we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Therefore, the type II error is at most
$2|\mathcal{B}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.7 Proof of Lemma 4
Proof.
We prove this lemma in three steps.
Step 1. sub-exponential norm for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
For the random variable $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$, applying (5) we obtain
$D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}=\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\big{(}\log(p^{A}_{n}(y_{n}))-\log(p^{B}_{n}(y_{n}))\big{)}\leq\sum_{y_{n}\in\mathcal{X}}-p^{A}_{n}(y_{n})\log(p^{B}_{n}(y_{n}))\leq-\log(\epsilon),$ (12)
where the first inequality holds since $\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log(p_{n}^{A}(y_{n}))\leq 0$, and the second inequality holds by Assumption 1.
For the random variable to be sub-exponential, by Definition 1, we need to find $t$ such that
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq 2.$
Applying (12), for the random variable to be sub-exponential with norm $t$, it is sufficient to satisfy
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq e^{\frac{-2\log(\epsilon)}{t}}\leq 2.$
Hence, $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ has sub-exponential norm $-4\log(\epsilon)$.
Step 2. Concentration bounds for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
Following the same steps as in the proof for Theorem 1, we conclude that $\sum_{n=1}^{N}[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Next, setting $t=c_{4}N$, we obtain
$\mathbb{P}\bigg{(}\big{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\big{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (13)
Step 3. Tail bound for cross-entropy
In this proof, for notation brevity we write $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$ instead of $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$, and $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\big{]}$ instead of $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}\big{]}$.
First, by Theorem 1 in Reeb and Wolf (2015), we have
$\Big{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}\leq\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K).$ (14)
Therefore, we have
$\Big{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}$ (15)
$=$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\bigg{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K)$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}$
$\geq$ $\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}\bigg{(}\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}-\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}\bigg{)}$
$\geq$ $\frac{1}{2N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{4N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]},$
where the first equality uses the definition of cross-entropy. The first inequality follows from the triangle inequality. The second inequality follows form (14). The third inequality follows from Assumption 3. The fourth inequality follows from QM-AM inequality. The last inequality follows from the difference of squares identity.
Hence, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\leq 2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}-\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq c_{4}\bigg{)}$
$\leq$ $2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]},$
where the first inequality follows from (15). The fourth inequality holds under Assumption 3, which requires $\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq 2\log^{2}(K)$ and also under the assumption in the statement of Lemma 4, which requires that $2c_{4}\leq\log^{2}(K)$. Finally, the last inequality follows from (13). This concludes the proof. ∎
A.8 Proof of Proposition 3
Proof.
Type I error occurs if a model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Applying triangle inequality, we have
$\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}.$
Hence, the type I error is upper bounded as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$ (16)
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}.$
From Lemma 4, we know that
$\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (17)
From Theorem 2 we know that
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (18)
Combining Equations (16-18), we conclude the type I error is upper bounded as
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if the model $A$ generates the text $\bf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t.$
We upper bound the probability of this event as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
where the inequality follows from Theorem 1. ∎
Appendix B Related work
Supervised detection methods. Supervised detection methods learn to differentiate between human-written and LLM-generated text through labeled examples. For example, GPT2 Detector (Solaiman et al. 2019) fine-tunes RoBERTa on the output of GPT2, and ChatGPT Detector (Guo et al. 2023) fine-tunes RoBERTa on the HC3 dataset in Guo et al. 2023. Many other supervised detection methods also exist, including those that leverage neural representations (e.g., Bakhtin et al. 2019; Uchendu et al. 2020; Fagni et al. 2021) and those that leverage bag-of-words (Fagni et al. 2021). Although some supervised detection methods show high performance, it has been observed that they tend to overfit their training domains and source models (Bakhtin et al. 2019; Uchendu et al. 2020). A key concern with neural-based methods (e.g., feature-based classifiers) is their poor robustness, e.g., against ambiguous semantics (Schaaff et al. (2023)) and their limited ability to detect LLM-generated misinformation (Schuster et al. (2020)). Recent efforts have been enhancing training methodology (e.g., Kumarage et al., 2023; Tu et al., 2023). However, these methods provide heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts. Additionally, there are limitations for the feature-based classifiers, including challenges in training models with the hype in developing new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns.
Zero-shot detection methods. Another stream of work attempts to distinguish machine-generated from human-written texts using statistics-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. Among those, white-box detection methods are closely related to our work. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures (Solaiman et al. 2019). Other measures include using GTLR based on rank-likelihood (Gehrmann et al. 2019), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. (2023), entropy and Kullback-Leibler (KL) divergence (Lavergne et al. 2008), and perplexity ( Vasilatos et al. 2023; Wang et al. 2023b). Mitchell et al. (2023) observes that machine-generated texts tend to lie in the local curvature of the log probability and proposes DetectGPT, which has demonstrated reliable performance but has some limitations. Du et al. analyzes the limitations of DetectGPT, highlights DetectGPT’s need for computing many perturbations, making it a computationally intensive algorithm, and proposes targeted masking strategies (rather than random masking) to improve DetectGPT’s performance. Hans et al. (2024) proposes Binoculars that use a ratio of perplexity measurement and cross-perplexity, a notion of how surprising the next token predictions of one model are to another model. Like supervised detection methods, most statistics-based detection methods provide only heuristics.
Watermarking. One detection approach is to record (Krishna et al., 2024) or watermark (Kirchenbauer et al., 2023) all generated text. The method proposed in Kirchenbauer et al. (2023) takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. Note that watermarking requires the cooperation of the generating party/LLM owner to implement the green listing-red listing algorithm. For a survey on watermarking methods, we refer readers to Amrit and Singh (2022).
Possibility of detection and robustness to attacks. Robust detection methods are being developed, and increasingly sophisticated evasion methods are being devised to circumvent these detectors, creating an ongoing contest between detection and evasion. The evasion of detection methods can be through (i) prompt engineering by general users who change the writing style of LLM-generated text or (ii) through paraphrasing attacks by adversaries. Recent research examines the ability of different detection methods against the mentioned evasion methods, including, e.g., Sadasivan et al. 2023; Krishna et al. 2024; Zhang et al. 2024. As a theoretical work in this stream, Chakraborty et al. (2023) provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$. In binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR). For a survey on possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. (2023).
Authorship attribution. The literature on human-written text author attribution is extensive and, similar to the recent literature on LLM text detection and attribution, has seen developments over time, e.g., use of statistical hypothesis testing versus discriminative methods like support vector machines, neural networks, etc. In author attribution with only human authors, there is no access to the text generation "model," and the attribution is only based on the samples of texts written by different human authors. For surveys on the different methods for human author attribution, we refer readers to Juola et al. 2008; Stamatatos 2009; Koppel et al. 2009, and for a comparison of different methods, we refer readers to Grieve (2007).
With the advent of LLMs, human author attribution extends to (1) human-written vs. machine-generated detection, (2) machine-generated vs. another machine-generated attribution, and (3) attribution in text generated through LLM and human collaboration. These shifts necessitate transitioning from human-centric stylometry to detection frameworks tailored for LLMs. For a survey that connects the older literature on authorship attribution to modern problems in the era of LLMs, we refer readers to Huang et al. (2025).
Statistical tests on LLM-generated text. Recent efforts have provided statistical tests for detecting LLM-generated text and determining whether cloud implementation works the same as the reference model. For example, Li et al. leverages the watermark key sequence and next token probabilities (NPTs) to introduce statistical tests for detecting LLM-generated text when the detection approach is watermarking, and their method applies to short and long text. Gao et al. (2024) introduces a framework for detecting whether an API provider has modified a language model’s output distribution without informing users. The paper formalizes the problem as a two-sample statistical testing task, where users collect samples from an API and compare them to a reference model’s outputs. They propose a hypothesis testing method using Maximum Mean Discrepancy (MMD) with a string kernel, first proposed in Gretton et al. (2012), to detect whether the API serves the expected model or a modified version.