### v4

# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

Abstract
Verifying the provenance of content is crucial to the function of many organizations, e.g., educational institutions, social media platforms, firms, etc. This problem is becoming increasingly challenging as text generated by Large Language Models (LLMs) becomes almost indistinguishable from human-generated content. In addition, many institutions utilize in-house LLMs and want to ensure that external, non-sanctioned LLMs do not produce content within the institution. In this paper, we answer the following question: Given a piece of text, can we identify whether it was produced by a particular LLM or not? We model LLM-generated text as a sequential stochastic process with complete dependence on history. We then design zero-shot statistical tests to (i) distinguish between text generated by two different known sets of LLMs $\mathcal{A}$ (non-sanctioned) and $\mathcal{B}$ (in-house), and (ii) identify whether text was generated by a known LLM or generated by any unknown model, e.g., a human or some other language generation process. We prove that the type I and type II errors of our test decrease exponentially with the length of the text. To do so, we show that if $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. We then present experiments using LLMs with white-box access to support our theoretical results and empirically examine the robustness of our results to black-box settings and adversarial attacks. In the black-box setting, our method achieves an average true positive rate (TPR) of 82.5% at a fixed false positive rate (FPR) of 5%. Under adversarial perturbations, our minimum TPR is 48.6% at the same FPR threshold. Both results outperform all non-commercial baselines. Our work has a host of practical applications where determining the origin of a piece of text is important, and can also be useful in combating misinformation and ensuring compliance with emerging AI regulations. See link for code, data, and an online demo of the project.
1 Introduction
LLM text generation tools are widely employed to produce textual content in various domains, including news agencies and academia. As LLM content generation tools improve, accurately telling whether a text is LLM-generated becomes increasingly challenging. But why must we know whether a language model generated the text? First, LLMs allow people to deliver content they did not produce. Second, LLMs allow content proliferation, including mass production of content deliberately tailored to specific audiences that can influence public discussions, such as manipulating financial markets [46], political discussions [61], or consumers’ sentiments [29]. Therefore, from the policymakers’ perspective, providing methods that reliably detect LLM-generated text is crucial. Alongside policymakers, generative AI providers must also determine AI-generated text because retraining on AI-generated content can lead to model collapse [55]. To avoid model collapse, AI providers must reduce the volume of AI-generated content in their training sets, which requires identifying such content reliably. The significance of AI-generated content detection is highlighted in [27], which has recently proposed the Fourth Law of Robotics: A robot or AI must not deceive a human by impersonating a human being. Implementing the fourth law would require clear labeling of AI-generated content, technical standards for AI identification, and legal frameworks for enforcement.
In addition to identifying LLM-generated text, distinguishing between text generated by different LLMs is critical. First, identifying the source of harmful content allows the assignment of responsibility in case of non-compliance with regulations (23, 69). Second, the quality of outputs can vary between models, and employing the most appropriate model is essential to achieving optimal results. In line with that, educational organizations are creating their LLMs to ensure educational integrity, e.g., [64]. Students can use only the specialized assistant for their assignments, not ChatGPT or other LLMs. This requirement highlights the need for reliable tools to detect text generated by prohibited LLMs. Alongside policymakers, generative AI providers have incentives to detect the text generated by their own language models to gauge the uptake of their systems, which is a commercially important measure of performance.
To evaluate the possibility of detecting LLM-generated text by (human) linguistics experts, [7] designed an experiment and reported an identification rate of only 38.9%. Since humans, even experts, struggle to detect LLM-generated text, researchers are designing detection methods to identify signs of LLM-generated text that are difficult to recognize for humans. One detection approach is to train neural networks on labeled training data from LLM and human classes (e.g., 3, 30, 63, 12, 52, 22, 66). Neural-based methods have limitations. First, they require training a separate (from the source model) classifier, which is challenging, especially considering the large and growing number of LLMs and the wide variety of topics and writing styles. Also, they require collecting datasets of human passages, which raises concerns, such as privacy, associated with training models on human data. Furthermore, [41] notes that because detectors are often evaluated on simple datasets, their performance on out-of-domain samples is typically abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using neural-based detectors.
Another detection method is watermarking [33], in which an LLM embeds hidden indistinguishable to human signals into the text during generation. Watermarking relies on cooperation from the AI company or the LLM owner, and current regulations cannot force companies to watermark their generated text [45]. The described limitations motivate the need for models that do not require training on human data or cooperation from the LLM owner. One such method is zero-shot training (see, e.g., the work of 42). The zero-shot detection studies leverage statistical properties of LLM-generated text for identification, including likelihood curvature [42], log-likelihood [56], rank-likelihood [16], log-likelihood ratio ranking [58], entropy and Kullback-Leibler (KL) divergence [38], and perplexity [65, 68, 24]. The existing detection methods are heuristic-based and lack formal guarantees.
To address this limitation, we provide a theoretical framework for determining whether a given finite length (possibly short) string was produced by LLM $A$ or another source $B$. We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single finite length string of text, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We first design tests to distinguish between two disjoint sets of known LLMs. For this purpose, we assume that we have white-box access to the models in the hypothesis test. In particular, we design composite tests that determine whether the text is generated by a model that belongs to a set of models $\mathcal{A}$ or to a disjoint set of models $\mathcal{B}$. We also study the case where we do not have white-box access to all models in the hypothesis set (for example, a human wrote the text) and design a composite statistical test to identify whether the text is generated by the model $A$ or not. To derive the type I and type II errors for our tests, we show that if $B$ generates the text, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$.
Contributions. We contribute to the literature on zero-shot statistical tests by developing the first statistical test with theoretical guarantees that identifies whether a finite length (possibly short) text was generated by a known LLM or not. Additionally, we are the first to provide theoretically guaranteed statistical tests that identify whether a finite (possibly short) text was generated by an LLM from an in-house set of known LLMs or a non-sanctioned set of known LLMs. This problem is becoming crucial with the development of specialized LLMs. We also prove that our tests’ type I and type II errors decrease exponentially as the text length increases.
To establish our theoretical results, we consider a sequence of discrete random variables $Z_{n}$ generated under probability laws $p_{n}^{B}(\cdot)$. We propose and prove an exponential decay concentration inequality that bounds the tail probability of the difference between the log-perplexity of this sequence evaluated under another sequence of probability laws $p_{n}^{A}(\cdot)$, and the average cross-entropy of $p_{n}^{B}(\cdot)$ and $p_{n}^{A}(\cdot)$, over a finite alphabet. The special case where (i) the sequence $Z_{n}$ is independent and identically distributed (iid), and (ii) $p_{n}^{B}(.)$ and $p_{n}^{A}(.)$ are the same is studied in [72]. However, LLMs generate text sequences dependent on the previously generated tokens. Therefore, we generalize the results in [72], relaxing (i) and (ii). Interestingly, we address one of the future research directions mentioned by [72], which is considering a sequence of non-independent random variables.
This paper is organized as follows: In Section 2, we present our mathematical model. In Section 3, we provide the concentration bounds. In Section 4, we present our statistical tests and derive the upper bound on their type I and type II errors. In Section 5, we show the results of our experiments. We conclude the paper in Section 6. Appendix A presents the proofs, Appendix B presents additional experiments and experiment details, and Appendix C provides a literature review.
2 Model and background definitions
2.1 Model
Let $M$ be a generative model described by $\textbf{Y}=m(\textbf{X})$, where X denotes user prompt and the output denoted by Y consists of a string of tokens $\textbf{Y}=[Y_{1},Y_{2},\dots,Y_{N},\dots]$. Each token is chosen from a finite vocabulary set, i.e., $Y_{n}\in\mathcal{X}$ with size $K:=|\mathcal{X}|$. Practical implementations of LLMs specify the probability distribution iteratively, e.g., 50. The model first draws a random value for the first token, say $Y_{1}=y_{1}$ by sampling from the distribution $p^{M}(Y_{1}|\textbf{X})$, and then for each token $n\in[2,N]$, the model sequentially determines a distribution for the token given prompt X and all the randomly chosen values $y_{1},y_{2},\dots,y_{n-1}$. So, we define a sequence of probability distributions $p^{M}(\textbf{Y}_{N}|\textbf{X})$ over $\textbf{Y}_{N}\in\mathcal{X}^{N}$ where $\textbf{Y}_{N}=[Y_{1},Y_{2},\dots,Y_{N}]$ is a substring of Y consisting of the first $N$ tokens. The sequence of probability distributions is determined as
$P^{M}(\textbf{Y}_{N}|\textbf{X})=\prod_{n=1}^{N}p^{M}_{n}(\textbf{Y}_{n}),\quad\text{where}\quad p^{M}_{n}(\textbf{Y}_{n})=P^{M}(Y_{n}|Y_{1},Y_{2},\dots,Y_{n-1},\textbf{X})$
Note that Equation 1 is an application of the Bayes’ rule and holds for any generative model (regardless of whether tokens $Y_{n}$ are sequentially generated). While equation 1 holds for all generative models, because conditional distributions $p_{n}(y)$ are in general not easily accessible, we apply the rule for sequential models. For our theoretical derivations, we follow the literature on white-box detection, and assume that we know the probability law $p^{M}_{n}(\textbf{Y}_{n})$ for any given sequence $\textbf{Y}_{n}$. See, for example, Mitchell et al. 42, Gehrmann et al. 16  [^1] Open-weight models, including LLAMA, DeepSeek, Mistral, Falcon, and MPT, provide access to the model, which includes access to the (conditional) log probabilities. Additionally, multiple closed-weight models also provide log probabilities through API commands. For example, OpenAI provides log probabilities through the Chat Completion API Command. Also, the language models supported by Together AI provide log probabilities by adding ”logprobs: 1” to the request.. For experiments with no white-box access to the log-probabilities, please refer to Section 5.3.
2.2 Background definitions
Consider an evaluator model $A$ and a text generator model $B$.
Perplexity: The perplexity $p^{A}(Y_{N})$ of a finite length text string $\textbf{Y}_{N}=[Y_{1},Y_{2},\ldots,Y_{N}]$ with respect to an evaluator model $A$ is defined as the per-token inverse likelihood of the string Y. Formally, perplexity with respect to model $A$ is $p^{A}(\textbf{Y}_{N})=\left(\prod_{n=1}^{N}p^{A}_{n}(Y_{n})\right)^{-\frac{1}{N}}$, and the log-perplexity is
$l_{A}(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\log(p^{A}_{n}(y_{n}))$
Intuitively, log-perplexity measures how "surprising" a string is to an evaluator language model.
Cross-entropy: Suppose model $B$ generates a string $\mathbf{Y}$. The average cross-entropy of model $B$ and the evaluator model $A$ over sub-string $\mathbf{Y}_{N}$ is
$h_{N}(B,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$
Intuitively, the average cross-entropy measures how surprising the token predictions of $B$ are when observed by $A$. If model $A$ is both the string’s generator and evaluator, then (average) cross-entropy and (average) entropy are equivalent. Our consideration of white-box detection yields that we know the probabilities $p^{A}_{n}(y_{n})$, and therefore, we can compute the average entropy $h_{N}(A,A)$.
Random Variables. We define the random variable $Z_{n}^{A}:=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and the zero-mean random variable $X_{n}^{A}:=Z_{n}^{A}-\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]$. Additionally, we define a random variable $S_{N}^{A}:=\sum_{i=1}^{N}X^{A}_{i}$. Note that we only define the random variable $X_{n}^{A}$ to obtain our concentration inequalities, and our statistical tests in Sections 4 do not require calculating $X_{n}^{A}$. The tests only require access to the evaluator model $A$ to compute $Z_{n}^{A}$ and $\mathbb{E}_{p^{A}_{n}}[Z_{n}^{A}]$. Given our white-box assumption, we have access to and can compute $Z_{n}^{A}$ and $\mathbb{E}_{p^{A}_{n}}[Z_{n}^{A}]$.
Consider a string $\bf{Y}$ generated by model $B$, and we want to evaluate the string using model $A$. So, $\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]$ is
where $H(p^{B}_{n},p^{A}_{n})$ is the cross-entropy between distributions $p^{B}_{n}(.)$ and $p^{A}_{n}(.)$. Note that $\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$ is not necessarily finite. From equation (4), $\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$ is infinite if there exists $y_{n}\in\mathcal{X}$ such that $p_{n}^{B}(y_{n})>0$ and $p_{n}^{A}(y_{n})=0$. If such a token $y_{n}$ appears in $\textbf{Y}_{N}$, then model $A$ assigns zero probability to the sequence, and we can conclude with certainty (i.e., probability 1) that model $A$ did not generate the string. Hence, the detection problem becomes trivial. Otherwise, i.e., $\textbf{Y}_{N}$ does not include any $y_{n}$ with $p^{A}_{n}(y_{n})=0$, we calculate $\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$ by applying conditional expectation as
[…]
where for $p^{A}_{n}(y_{n})=0$, we have $p^{\tilde{B}}_{n}=0$ and for $p^{A}_{n}(y_{n})>0$, we have $p^{\tilde{B}}_{n}=p^{B}_{n}(y_{n})/\sum_{y_{k}:p^{A}_{n}(y_{k})>0}p^{B}_{n}(y_{k})$. Hence, the detection problem for model pairs $A$ and $B$ reduces to that of models $A$ and $\tilde{B}$. The modified pair of distributions $A,\tilde{B}$ satisfy the condition that if $p^{A}_{n}(y_{n})=0$, then $p^{\tilde{B}}_{n}(y_{n})=0$ for all $y_{n}$, which yields $\mathbb{E}_{p_{n}^{\tilde{B}}}[Z_{n}^{A}]$ is finite. Without loss of generality, we exclude the trivial case and only focus on pairs of models for which $\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$ is finite.
3 Concentration inequalities
In this section, we provide concentration inequalities to show that if the string is generated by a model $B$ and evaluated by a model $A$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$ converges to the average cross-entropy of the string under $B$ and $A$, $h_{N}(B,A)$, with a high probability. If the text is generated and evaluated by model $A$, then $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$ converges to the average entropy of the string under $A$ with high probability. In Appendix A.1, we provide the proofs for the results in this section. These concentration bounds are the backbones of the statistical tests we design in section 4. For tractability, we make a parametric assumption on the probability laws $p^{A}_{k}(.)$ and $p^{B}_{k}(.)$.
Assumption 1.
We assume that there exists $\epsilon>0$ such that $p^{A}_{n}(y_{k}),p^{B}_{n}(y_{k})\notin(0,\epsilon)$.
Assumption 1 implies that models $A$ and $B$ either do not associate any probability to a token $y\in\mathcal{X}$, or they assign a probability of at least $\epsilon$. Our theoretical results depend only on $\log(\epsilon)$. Hence, our theoretical bounds primarily rely on a constant shift in the logarithmic scale. Note that Assumption 1 is not restrictive and is aligned with practice, as computers only allow for a limited range of representable numbers due to finite precision in floating-point arithmetic. Very small probabilities are either rounded to zero or set to a minimum threshold to maintain numerical stability in computations (18). The first outcome of assumption 1 is that
With $\mathbb{E}_{p_{n}^{B}}\big{[}Z_{n}^{A}\big{]}<\infty$, we define the zero-mean random variable $X_{n}^{A}=Z_{n}^{A}-\mathbb{E}_{p_{n}^{B}}[Z_{n}^{A}]$. Theorem 1 presents our concentration inequality, which we apply to design our statistical tests.
Theorem 1.
(a) If model $A$ is the same as model $B$, there exists a constant $c_{1}>0$ independent of the evaluator model $A$ such that for any $t>0$ we have
$\mathbb{P}\bigg{(}\frac{1}{N}\left|\sum_{n=1}^{N}X_{n}^{A}\right|\geq t\bigg{)}\leq 2\exp\left[-\frac{Nt}{c_{1}\log(K)}\min\left(1,\frac{t}{c_{1}\log(K)}\right)\right].$
(b) If model $A$ is not model $B$, there exists a constant $c_{3}>0$ independent of models $A$ and $B$ such that for any $t>0$ we have
[…]
Proof.
For proof of part (a), see Section A.1.3 and for proof of part (b), see Section A.1.4. ∎
Interpretation. Theorem 1 - (a) states that if the string is generated by the same model as the evaluator model $A$, then the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. Theorem 1 - (b) states that if model $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under the evaluator model $A$ converges to the average cross-entropy of the string under $B$ and $A$.
Theorem 1 implies that for any evaluator model $A$, $\sum_{n=1}^{N}Z_{n}^{A}$ converges to $h_{N}(B,A)$. As such, for $A=B$, $\sum_{n=1}^{N}Z_{n}^{A}$ converges to $h_{N}(A,A)$. Hence, our main idea in designing a statistical test is to check if $\sum_{n=1}^{N}Z_{n}^{A}$ has converged to $h_{N}(A,A)$. Our next section formalizes this idea.
4 Statistical test
We design our statistical tests using the result in Theorem 1 and then evaluate type I (false positive) and type II (false negative) errors. In Section 4.1, we design a composite statistical test that determines whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or belongs to a disjoint set of models $\mathcal{B}$. In Section 4.2, we study the case where we do not have access to the log-probabilities for the generator model (e.g., text is written by an unknown LLM, a human, etc.) and design a composite test to identify whether the model $A$ generated the text or not. In Appendix A.2, we provide the proofs for the results in this section.
4.1 Attribution among multiple LLMs
Given a string $\mathbf{Y}_{N}$, we design a statistical test to detect whether one of the models in $\mathcal{A}=\{A_{1},\ldots,A_{p}\}$ (non-sanctioned) or one of the models in =$\mathcal{B}=\{B_{1},\ldots,B_{q}\}$ (in-house) generated the text. The null hypothesis $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is generated by a model in $\mathcal{B}$. We first calculate the random variables $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$, and sum over $Z^{M}_{n}=:-\log\big{(}p^{M}_{n}(Y_{n})\big{)}$ for all models $M\in\mathcal{A}\cup\mathcal{B}$. Our test rejects the null $\mathbf{H}_{0}$ if for some $A_{i}\in\mathcal{A}$, we have
Type I error occurs when the test incorrectly concludes that the text is generated by one of the models in $\mathcal{A}$, and Type II error occurs when the test fails to identify that the text is generated by one of the models in $\mathcal{A}$. Our theoretical bounds require that the two models do not impose the same probability distribution on the string. Hence, we introduce Assumption 2 as a minimum distance between log-probabilities for the two models.
Assumption 2.
(minimum difference-MM). We assume that if the generative and evaluator models are different, for an arbitrarily small $\epsilon_{1}>0$, we have
$\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\geq\epsilon_{1}.$
Assumption 2 ensures that the two models satisfy a minimum distance in terms of their KL divergence over the generated text  [^2] Using KL divergence as a measure for comparing the learnt distribution from a text generation model to the distribution of another text generation model or the distribution of human-written text is common, e.g., see 48 and the references within.. Note that KL divergence, by definition, is a non-negative value that demonstrates the distance between the two distributions over the next word for the two models. Our results show that the type I and type II errors for our statistical test are approximately $\exp\big{(}O(-N\epsilon_{1})\big{)}$, which indicates that even for small values of $\epsilon_{1}$, which can converge to zero with the text length (for example, $\epsilon_{1}=O(N^{-1/2})$), our test provides exponentially small type I and type II errors in the text length.
Proposition 1.
If Assumptions 1 and 2 hold, the type I error for our test is upper bounded by
and the type II error for our test is upper bounded by
with constants $c_{1}$, $c_{3}$, and $\epsilon$ as introduced in Theorem 1.
Proof.
See Section A.2.1. ∎
Interpretation. Proposition 1 demonstrates that the type I and type II errors of our composite test decrease exponentially in the text length.
4.2 Model $A$ or not model $A$?
Given a string $\mathbf{Y}_{N}$, for arbitrary constants $t$, we design a statistical test to detect whether the evaluator model $A$ generated the text. In hypothesis testing, the null hypothesis reflects the default/conservative assumption. So, if we accuse someone of having used LLM $A$, they are innocent until proven guilty. Accordingly, the null $\mathbf{H}_{0}$ is that the text $\mathbf{Y}_{N}$ is not generated by the evaluator model $A$; it is generated by any model $B$ that we do not have white-box access to, which could be a human or some other language generation process [^3] Recall from section 2.1 that an application of the Bayes’ rule yields that any text generation process corresponds to some sequential model $B$ regardless of whether tokens are sequentially generated. Our test does not require any information about $B$.. The alternative $A$ is that it is generated by $A$. We first calculate the random variable $Z_{n}^{A}=:-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and then we calculate $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}$. Our test rejects the null if
[…]
Type I error occurs when the test mistakenly concludes that the text is generated by the evaluator model $A$, and Type II error happens when the test incorrectly concludes that the text is not written by $A$. To derive type I and type II error controls, we impose Assumption 3. Note that if the two distributions are identical, no test can distinguish them, and, therefore, some assumption about distributions is always needed for any test to work.
Assumption 3.
(minimum difference-HM). We assume that if the generative and evaluator models are different, then for some positive $\epsilon_{2}>0$
We make two points about this assumption: First, note that $h_{N}(A,A)$ is always upper bounded by $\log(|\mathcal{X}|)$. Second, an objective in designing LLMs is to reduce the expected log-perplexity, which is equivalent to the average entropy $h_{N}(A,A)$, e.g., for language models similar to GPT-3, $h_{N}(A,A)$ is less than 4 (47). Our theoretical results only require that $|h_{N}(B,A)-h_{N}(A,A)|>\epsilon_{2}$ for an arbitrary $\epsilon_{2}>0$, and Assumption 3 is a sufficient condition that ensures that an $\epsilon_{2}>0$ exists.
Proposition 2.
For any $t\geq 0$, the type II error for our test is upper bounded by
with $c_{1}$ as introduced in Theorem 1. Also, if Assumptions 1 and 3 hold, then for any positive numbers $t<\epsilon_{2}/2$, and for $c_{3}$ as introduced in Theorem 1, the type I error of our test is upper bounded by
Proof.
See Section A.2.2. ∎
Interpretation. Proposition 2 demonstrates that the type I and type II errors of our test decrease exponentially in the text length. We note that our theoretical results also apply to a one-sided test that rejects the null hypothesis when $\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\leq t.$
5 Experiments
This section presents three sets of experiments: Subsection 5.1 examines Proposition 1, and Subsection 5.2 examines Proposition 2 where we use human text to represent the null hypothesis (not LLM $A$). Subsection 5.3, using a dataset with approximately 6M texts, empirically examines the robustness of Proposition 2 to (i) relaxing white-box access, and also (ii) to an adversary manipulating texts that affects both type I and type II errors.
For experiments on the convergence of the log-perplexity to the average string entropy, please see Appendix 5.
5.1 LLM attribution experiments
For our experiments in this section, we use two datasets that cover a variety of domains and LLM use-cases. We use news articles from the XSum dataset [44] to represent fake news detection, and Wikipedia paragraphs from SQuAD contexts [51] to represent LLM-generated academic essays. For details about datasets, please refer to Appendix B.2. To create LLM text, we first randomly sample 500 texts from each dataset. Then, for each sample, we prompt the first 30 tokens to each LLM and ask it to generate a 300-token text. The language models we use for generating the 300-token texts are LLaMA2 (7B) and LLaMA3 (8B) from the same family, Falcon (7B), and Qwen3 (32B) as a large model. We have white-box access to all models, which aligns with the Proposition 1’s requirement. In attribution experiments for each LLM, the evaluator LLM is the set $\mathcal{A}$, and the other three LLMs are the set $\mathcal{B}$. The null hypothesis is that the text is generated by one of the models in $\mathcal{B}$. Type I error occurs when the test incorrectly concludes that the text is generated by the evaluator model. Type II error occurs when the test fails to notice that the evaluator model generated the text.
Results in Table 1 demonstrate the high accuracy of our framework in the LLM attribution task. The LLM in the headline of each column stands for the evaluator LLM.
XSum (prefix = 300) SQuAD (prefix = 300)
Error Type LLaMA2 LLaMA3 Falcon Qwen3 LLaMA2 LLaMA3 Falcon Qwen3
Type I Error 0.02 0.00 0.02 0.00 0.02 0.00 0.00 0.02
Type II Error 0.00 0.14 0.00 0.00 0.00 0.10 0.04 0.00
Table 1: Type I and Type II error rates (rounded to 2 digits) for LLM attribution on XSum and SQuAD datasets.
5.2 Detection experiments
Setup & Datasets. To examine Proposition 2, we use human-written text to represent the null hypothesis (not LLM $A$). For human-written text, in addition to the Xsum and SQuAD datasets, we use prompted stories from the Reddit WritingPrompts dataset [13] to represent human creative writing. For details about datasets, please refer to Appendix B.2. We randomly sample 500 texts from each dataset, and prompt the first 30 tokens of each text to the LLM to create an LLM-generated version. The language models we examine are LLAMA 3 (8B), GPT-NEOX Erebus (20B), and QWEN (32B). We have white-box access to all models, which aligns with the Proposition 2’s requirement.
Test & Metrics. For this section, since the first 30 tokens are generated by a human for both null and alternative, we use the one-sided version of our test that removes the impact of the differences between cross-entropy and log perplexity caused by the initial 30 tokens. For a detailed explanation, please refer to Appendix B.3. Here, True Positive Rate (TPR) is the rate of correctly spotting that the evaluator model generated the text, and False Positive Rate (FPR) is the rate of incorrectly flagging human text as generated by the evaluator model. Note that FPR is equivalent to Type I error, and TPR is equivalent to 1-Type II error. Unlike Proposition 1, which is threshold-independent, Proposition 2 relies on a fixed threshold $t$ for rejecting the null hypothesis. Varying this threshold affects both the FPR and the TPR. Since different detectors use distinct methodologies and require their own thresholds, it is essential to account for the impact of these thresholds when evaluating the test performance. To address this, the literature typically adopts two approaches: (1) evaluating the TPR at a specified low FPR, e.g., [24, 11], or (2) using the Area Under the Receiver Operating Characteristic (AUROC) curve as a summary measure of performance, e.g., [66, 42]. Accordingly, we use two metrics. Our first metric is TPR, while FPR is held less than or equal to 1%. Note that the results of experiments in the literature, e.g., in [11], suggest that the FPR / type I error for current detectors can be dangerously high, which makes them unfit for regulatory applications. Our experiments using a TPR at a low FPR metric show the validity of our theoretical results and our method’s potential for regulatory application. Our second metric is AUROC, which plots TPR against FPR at all possible thresholds.
Benchmarks. We compare the performance of our method with various zero-shot methods. In choosing the benchmarks, we follow the literature, e.g., [42]. The log-probability-based benchmarks use the source model’s average token-wise log-probabilities. The (log-)rank-based benchmarks use the average observed (log-)rank of the tokens in the passage based on the model’s conditional distributions. The entropy-based benchmarks hypothesize that LLM texts are less-surprising for the model, leading to lower entropy predictive distributions. We also use Binocular [24] which has shown to substantially outperform all the zero shot detectors proposed in the literature in terms of TPR at low FPR.
Detection results. Table 2 shows TPR, while FPR is held less than or equal to 1%. The results show the higher performance of our method in fact-based datasets (e.g., SQuAD), where the human and LLM texts are less different, resulting in lower performance of the existing detection methods. Even when the FPR is held low, the human and LLM texts are more distinguishable in datasets with more creative content (e.g., WritingPrompts). Also, we observe that smaller language models (e.g., LLAMA 3 (8B) are easier to detect even with current detectors, and the maximum gap between ours and the second-best methods is in the SQuAD dataset with QWEN (32B) as the evaluator. Results for the AUROC are in Appendix B.1.
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.89 0.02 0.99* 0.01 0.05 0.01 1.00 0.46 1.00
Rank 0.22 0.06 0.20 0.07 0.18 0.01 0.49 0.39 0.43
LogRank 0.99 0.05 0.99* 0.02 0.06 0.01 1.00 0.54 1.00
Entropy 0.02 0.09 0.01 0.01 0.04 0.00 0.00 0.00 0.00
Binoculars 0.99 0.89* 0.99* 0.96 0.86* 0.26* 0.97* 0.99 0.99
Ours 0.97* 1.00 1.00 0.96 0.92 0.72 0.90 0.99 1.00
Diff (zero-shot) -0.02 0.11 0.01 0.00 0.06 0.46 -0.10 0.00 0.00
Table 2: TPR at FPR=0.01. The best TPR values are in bold, and the second-best values are marked with an asterisk (*). The row Diff(zero-shot) shows our TPR improvement over the strongest zero-shot baseline methods.
5.3 Robustness experiments
Setup, Datasets & Metrics. In our theoretical results, we assumed white-box access to the generator models. In this section, we empirically evaluate the robustness of the statistical test in Proposition 2 under the black-box setting. As in section 5.2, we consider human text as representing the null hypothesis and text generated by a language model as representing the alternative hypothesis. We do not have access to the specific model that generated the text in the alternative set.
Instead, we leverage a standard assumption in the literature that language models are more similar to each other than to human text to design our test (e.g., see section 4 in 48). In the context of Proposition 2, this implies that the cross-entropy between human text and a language model is higher than the cross-entropy between two language models. We use the scoring model tiiuae/falcon-rw-1B to compute the log-perplexity and average entropy of text samples, and then apply the test as defined in Proposition 2. This evaluation setup is consistent with prior work on black-box detection (e.g., 42, 11). Following [11], we use TPR at 5% FPR as our evaluation metric.
We use the RAID dataset [11] that consists of 14,971 human-written texts from public pre-2022 datasets from 8 domains of abstracts, books, news, poetry, recipes, Reddit, reviews, and Wikipedia. For each human text, [11] creates a corresponding generation prompt using a template such as "Write a recipe for title". Then, they generate one output for each of the 11 models, 4 decoding strategies, and 11 adversarial attacks. The 11 models are: open-source chat models Llama-c, Mistral-c, and MPT-c; open-source non-chat models Mistral, MPT, and GPT2; closed-sourced chat models c-GPT, GPT4, and Cohere; and closed-sourced non-chat models Cohere and GPT3. They consider 4 decoding strategies: (i) greedy, which picks the highest probability token at each generation or sampling, which samples according to a generation probability $p_{n}^{B}(.)$, and (ii) repetition penalty or no repetition penalty. The repetition penalty evades detectors by reducing the likelihood of generating previously used tokens. This is achieved by down-weighting their probabilities by a factor $\theta\in\{1,1.2\}$, resulting in less repetitive output. After generating each text, they apply adversarial attacks that manipulate text on all samples, including human text and the $11\times 4$ LLM generations. These manipulations induce type I and type II errors in the statistical test. The nine query-free attacks are: Alternative Spelling (AS), Article Deletion (AD), Insert Paragraphs (IP), Number Swap (NS), Misspelling (MS), Synonym Swap (SS), Upper Lower Swap (ULS), Paraphrasing (PP), and Whitespace Addition (WSA). For details about decoding strategies and adversarial attacks, please refer to Appendix B.4. The described process yields a total of 134,739 human texts and 5,928,516 LLM texts. Given the cost of more than \$10,000 for re-running the experiment for the performance of benchmark detectors on approximately 6M texts, we use the results of [11]’s experiments on the other detectors and run the experiment for our own detector. For details about hardware and computation costs, refer to Appendix B.6.
Benchmarks. We compare our performance with multiple neural, zero-shot, and commercial detectors. The neural benchmarks are RoBERTa-B (GPT-2), RoBERTa-L (GPT-2), RoBERTa-B (ChatGPT), and RADAR. The zero-shot benchmarks are GLTR, Fast DetectGPT, Binoculars, and LLMDet. The commercial benchmarks are GPTZero, Originality, Winston, and ZeroGPT. For details about detectors, please refer to Appendix B.5. Recall that in this section, we do not have white-box access, and we use tiiuae/falcon-rw-1B as our scoring model.
Black-box detection results. Results in Table 3 demonstrate high performance of our detector in black-box detection with and without sampling and repetition penalty. For example, while other detectors have low TPR when detecting open-source non-chat models with sampling and repetition penalty, we achieve a TPR of 73.3%. Furthermore, we observe that, unlike all other zero-shot and neural-based methods, the performance of our framework never falls below the low threshold, with our lowest performance being 54.0%. The average performance of our detector in black-box detection is 82.9%, which (except for Originality, a commercial detector with TPR=85.2%) is higher than the average performance of other detectors, with the second highest performance being Binoculars with TPR=80.6 %.
Open-Source Closed-Source
Chat Models Non-Chat Models Chat Models Non-Chat Models
(llama-c, mistral-c, mpt-c) (mistral, mpt, gpt2) (c-gpt, gpt4, cohere) (cohere, gpt3)
Dec. Strategy greedy sampling greedy sampling greedy sampling greedy sampling
Rep. Penalty? $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
R-B GPT2 84.1 52.3 77.9 26.2 98.6 44.1 60.5 35.4 70.9 41.7 65.1 52.5
R-L GPT2 79.7 41.1 71.4 19.5 98.5 43.0 67.2 53.4 61.4 34.7 61.1 48.6
R-B CGPT 80.2 63.3 75.0 39.3 53.3 26.4 14.9 1.7 59.1 38.1 46.5 39.0
RADAR 88.8 77.4 85.6 66.4 91.8 63.8 48.3 31.8 81.6 75.3 72.2 67.7
GPTZero 98.8 93.7 98.4 82.5 74.7 34.6 9.4 4.8 92.3 88.5 60.6 53.4
Originality 98.6 86.3 97.7 72.5 99.9 64.1 89.0 51.2 96.8 89.0 91.7 85.4
Winston 97.2 90.1 96.6 78.3 68.2 49.0 29.5 11.3 96.1 93.7 73.2 68.1
ZeroGPT(*) 95.4 80.7 90.5 54.9 85.1 57.2 16.0 0.3 92.1 65.8 83.4 72.7
GLTR 89.8 67.5 83.9 38.3 99.6 56.9 44.5 0.5 80.7 54.3 75.6 63.7
F-DetectGPT 98.6 74.5 96.2 40.5 97.8 56.1 79.7 0.6 96.0 74.1 93.8 86.3
LLMDet 55.5 30.2 47.5 16.5 74.8 27.0 38.4 3.7 35.8 18.5 40.0 32.9
Binoculars 99.9 86.6 99.7 60.6 99.9 62.3 72.4 0.6 99.2 92.1 99.0 95.0
Ours 99.4 86.4 97.9 62.2 90.8 84.0 54.0 73.3 96.4 70.2 94.4 86.3
Table 3: TPR at FPR=5%. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
Adversarial attacks results. Our detector maintains relatively high accuracy across multiple adversarial scenarios, with our lowest TPR at FPR=5% being 48.6%, which is higher than the lowest performance of all other zero-shot and neural-based detectors. Like other zero-shot detectors, our performance is reduced under Synonym Swap and Whitespace Addition, which is predictable because zero-shot methods often rely on fixed token distributions.
None AS AD IP NS MS SYN ULS WSA PP
RoB-B GPT2 59.1 55.6 37.1 56.9 55.9 43.8 71.5 18.8 45.2 68.9
RoB-L GPT2 56.7 52.4 33.2 55.1 51.7 39.5 79.4 19.3 40.1 72.9
RoB-B CGPT 44.8 43.3 38.0 5.2 44.3 42.1 39.6 31.7 0.1 49.2
RADAR 70.9 70.8 67.9 73.7 71.0 69.5 67.5 70.4 66.1 67.3
GPTZero 66.5 64.9 61.0 66.2 65.8 65.1 61.0 56.5 66.2 64.0
ZeroGPT 65.5 65.4 59.7 64.9 64.7 64.7 18.8 54.5 64.2 46.7
Originality 85.0 83.6 71.4 85.1 86.0 78.6 96.5 75.8 84.9 96.7
Winston 71.0 68.9 66.9 69.8 69.0 67.5 63.6 56.8 46.8 52.6
GLTR 62.6 61.2 52.1 61.4 59.9 59.8 31.2 48.1 45.8 47.2
F-DGPT 73.6 71.6 64.7 72.0 68.2 70.7 34.0 60.4 64.4 71.8
LLMDet 35.0 33.9 27.4 27.2 33.8 32.7 27.3 23.4 4.4 28.5
Binoculars 79.6 78.2 74.3 71.7 77.1 78.0 43.5 73.8 70.1 80.3
Ours 82.5 80.6 74.4 81.1 74.8 79.8 48.6 72.5 60.8 74.2
Table 4: TPR @ FPR=5%. Abbreviations are: AS: Alternative Spelling, AD: Article Deletion, IP: Insert Paragraphs, NS: Number Swap, MS: Misspelling, SYN: Synonym Swap, ULS: Upper Lower Swap, WSA: Whitespace Addition, PP: Paraphrase. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
6 Conclusion
We answer the following question: Given a piece of text, can we identify whether it was produced by a particular LLM or not? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single finite length string of text, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We first design composite tests that determine whether the text is generated by a model that belongs to a set of models $\mathcal{A}$ or to a disjoint set of models $\mathcal{B}$. We then design a composite statistical test to identify whether a text is generated by the model $A$ or not. In Propositions 1 and 2, we prove that our tests’ type I and type II errors decrease exponentially as the text length increases. As a prerequisite for designing our tests, in Theorem 1, we prove an exponential decay concentration inequality that bounds the tail probability of the difference between the log-perplexity of a sequence generated by $B$ and evaluated under $A$, and the average cross-entropy of $B$ and $A$, over a finite alphabet. The special case where (i) the sequence is independent and identically distributed (iid), and (ii) models $B$ and $A$ are the same, is studied in [72]. However, LLMs generate text sequences that depend on the previously generated tokens. Therefore, our result in Theorem 1 generalizes the results in [72], relaxing (i) and (ii). Interestingly, we address one of the future research directions mentioned by [72], which is considering a sequence of non-independent random variables. Our numerical experiments in a white-box setting show the validity of our theoretical results. In the black-box setting, our method achieves an average TPR of 82.5% at a fixed FPR of 5%. Under adversarial perturbations, the minimum TPR is 48.6% at the same FPR threshold. Both results outperform all non-commercial baselines. We hope our work inspires more research on zero-shot LLM-text detection with provable guarantees.
Appendix A Appendix
A.1 Concentration inequality proofs
In this section, to improve readability, we first provide the prerequisites of sub-exponential random variables that are used in our proofs in Section A.1.1, followed by a review of notations in Section A.1.2. We provide the proof for Theorem 1 - part (a), i.e., where the generator model $B$ and the evaluator model $A$ are the same, in Section A.1.3. The starting paragraph in the proof provides a high-level overview, describing the proof steps, remarks, and lemmas used as part of the proof. In Section A.1.4, we provide the proof for Theorem 1 - part (b), i.e., where the generator model $B$ is not the same as the evaluator model $A$. Similar to the previous part, the starting paragraph of this section provides a proof overview.
A.1.1 Prerequisites of sub-exponential random variables from the literature
Before presenting the proofs for our main theorem, we need to clearly state the following definitions and the equivalency lemma for sub-exponential random variables from the literature.
Definition 1.
(sub-exponential norm). The sub-exponential norm of $X\in\mathbb{R}$ is
$\|X\|_{\psi_{1}}=\inf\bigg{\{}t>0:\mathbb{E}[e^{\frac{|X|}{t}}]\leq 2\bigg{\}}.$
If $\|X\|_{\psi_{1}}$ is finite, we say that $X$ is sub-exponential.
We also need the following equivalent (up to a constant factor) definition of a sub-exponential random variable from [67].
Definition 2.
(sub-exponential random variable). Centered random variable $X\in SE(\nu^{2},\alpha)$ with parameters $\nu,\alpha>0$ is sub-exponential if
$\mathbb{E}[e^{\lambda X}]\leq e^{\frac{\lambda^{2}\nu^{2}}{2}},\quad\forall\lambda:|\lambda|<\frac{1}{\alpha}.$
Next, we present a lemma from Vershynin [67] that demonstrates that the two definitions are equivalent up to a constant factor.
Lemma 1.
(SE properties ,67). Let $X$ be a random variable with $\mathbb{E}[X]=0$. Then, there exists a constant $c$ and constants $K_{4}$ and $K_{5}$ such that $K_{4}\leq cK_{5}$ and $K_{5}\leq cK_{4}$ and the following two properties are equivalent.
  • There exists a constant $K_{4}$, such that the MGF of $|X|$ is bounded, specifically
    $\mathbb{E}[e^{|X|/K_{4}}]\leq 2.$
  • There exists a constant $K_{5}$, such that the MGF of $X$ satisfies
    $\mathbb{E}[e^{\lambda X}]\leq e^{\frac{K_{5}^{2}\lambda^{2}}{2}}\quad\forall\lambda\quad s.t.\quad|\lambda|\leq\frac{1}{K_{5}}$
A.1.2 Review of notations
Here, we first review the notation in Section 2: X denotes user prompt and Y denotes the output that consists of a string of tokens $\textbf{Y}=[Y_{1},Y_{2},\dots,Y_{N},\dots]$. Each token is chosen from a finite vocabulary set, i.e., $Y_{n}\in\mathcal{X}$ with size $K:=|\mathcal{X}|$. We defined the random variable $Z_{n}^{A}:=-\log\big{(}p^{A}_{n}(Y_{n})\big{)}$, and the zero-mean random variable $X_{n}^{A}:=Z_{n}^{A}-\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]$. Additionally, we defined a random variable $S_{N}^{A}:=\sum_{i=1}^{N}X^{A}_{i}$.
A.1.3 Proof of Theorem 1 - part (a)
We first provide a high-level structure for our proof. Recall that in part (a), we focus on the case where the generator model $B$ is the same as the evaluator model $A$. In Remark 1, we show that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $\log K$, and therefore it is finite. In Lemma 2, we show that the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale. In Lemma 3, we find the sub-exponential norm for the martingale increments random variable $X_{n}^{A}$. As the last step, we show that martingale sum $\sum_{n=1}^{N}X_{n}^{A}$ is sub-exponential with parameters $(\nu^{2}N,\nu)$, with $\nu=c_{1}\log K$. Finally, we apply the tail bound for sub-exponential random variables to bound the probability of $|\sum_{n=1}^{N}X_{n}^{A}|$ being larger than a constant $t$. A challenge in applying the common concentration bounds for martingales is that martingale increments are not necessarily bounded. We overcome this issue by showing that the martingale differences, while not bounded, admit a light tail. In particular, we show that the martingale differences are sub-exponential.
Remark 1.
$\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]=-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})\leq\log K.$
Proof.
Concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ yields that its maximizer is $p^{*A}_{n}(y_{n})=\frac{1}{K},\forall y_{n}\in\chi$. Hence,
$\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]=-\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log p_{n}^{A}(y_{n})$$\leq$ $\sum_{y_{n}\in\chi}p_{n}^{*A}(y_{n})\log p_{n}^{*A}(y_{n})$
$=$ $K.\frac{1}{K}.\log K=\log K.$
Now that we have shown $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]<\infty$, recall that we defined the zero-mean random variable $X_{n}^{A}=Z_{n}^{A}-\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$. Then, we defined $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$, which as we show in Lemma 2, forms a martingale.
Lemma 2.
If the evaluator model $A$ is the same as the generator model $B$ that generates the string $\bf{Y}_{N}$, then the random variable $S_{N}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale.
Proof.
$Z_{n}^{A}$ and as a result $\mathbb{E}_{p_{n}^{A}}(y_{n})$ are positive random variables. In Remark 1, we showed that $\mathbb{E}_{p_{n}^{A}}[Z_{n}]\leq\log K$. Therefore, we have $\mathbb{E}\big{[}|X_{n}^{A}|\big{]}\leq\log K$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}^{A}|S_{N}^{A}]=S_{N}^{A}$.
Property (i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}=\mathbb{E}\big{[}|X_{N+1}^{A}|\big{]}\leq\log K<\infty$.
Property (ii) is satisfied because the martingale increments $X_{n}^{A}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Lemma 3.
The sub-exponential norm for the random variable $X_{n}^{A}$ equals $2\log(K)$.
Proof.
For the random variable to be sub-exponential, by Definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}^{A}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\frac{1}{p^{A}_{n}(y_{n})^{1/t}}=e^{\frac{\mathbb{E}[Z_{n}]}{t}}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq 2.$ (6)
where the first inequality follows from the triangle inequality and that $Z_{n}$ and therefore $\mathbb{E}[Z_{n}]$ are positive. Note that $\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}$ is concave in $p^{A}_{n}(y_{n})$. Hence, it attains its maximum when we have $p^{A}_{n}(y_{i})=p^{A}_{n}(y_{j})\quad\forall y_{i},y_{j}\in\mathcal{X}$, which yields
$\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})^{\frac{t-1}{t}}\leq K(\frac{1}{K})^{\frac{t-1}{t}}=K^{\frac{1}{t}}.$ (7)
To analyze $e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}$, we first want to show that
$\mathbb{E}[Z_{n}^{A}]$ $=$ $-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$ (8)
$\leq$ $\log(K).$
To show that, note that by concavity of $-p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$, we can maximize $p_{n}(y_{n})\log p_{n}(y_{n})$ by equalizing all probabilities $p^{*}(y_{n})=\frac{1}{K};\forall n$. So, we have
$-\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log p^{A}_{n}(y_{n})$$\leq$ $-\sum_{y_{n}\in\mathcal{X}}p^{*}(y_{n})\log p^{*}(y_{n})$
$=$ $-K\Big{[}\frac{1}{K}\log\frac{1}{K}\Big{]}=\log K.$
Combining equations (6)-(8) yields that for $X_{n}$ to be sub-exponential with norm $t$, it must satisfy
$K^{1/t}K^{1/t}\leq 2$
Hence, $X_{n}^{A}$ has sub-exponential norm $\|X\|_{\psi_{1}}=2\log(K)$. ∎
Proof of Theorem 1 part (a).
Proof.
Next, note that by Lemma 3 and Lemma 1, there exists $c_{1}>0$ s.t. for $\nu=c_{1}\log K$, and for any $n\in\mathbb{N}$,
$\mathbb{E}[e^{\lambda X_{n}^{A}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (9)
By Definition 2, we have $X_{n}\in SE(\nu^{2},\nu)$.
Our next step is to show that $\sum_{n=1}^{N}X_{n}$ is SE with parameters $(\nu^{2}N,\nu)$. To realize that, observe that
$\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n}X_{k}^{A}\right)}\right]=\mathbb{E}\left[e^{\lambda\left(\sum_{k=1}^{n-1}X_{k}^{A}\right)}\mathbb{E}\left[e^{\lambda X_{n}^{A}}\mid\mathbf{Y}_{n-1}\right]\right]\leq\mathbb{E}\left[e^{\lambda\sum_{k=1}^{n-1}X_{k}^{A}}\right]e^{\frac{\lambda^{2}\nu^{2}}{2}}\leq e^{\frac{\lambda^{2}N\nu^{2}}{2}},$
where the first equation follows from the iterated law of expectation, and the first inequality follows from Equation (9).
Finally, from Theorem 5.2 in 2, if $S\in SE(\nu^{2},\alpha)$ is a sub-exponential random variable, then
$\mathbb{P}(|S-\mathbb{E}[S]|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{\nu^{2}},\frac{t_{1}}{\alpha}\bigg{)}\bigg{)}.$ (10)
Substituting for the zero-mean random variable $\sum_{n=1}^{N}X_{n}$ that is sub-exponential with parameters $(\nu^{2}N,\nu)$, with $\nu=c_{1}\log K$, we obtain
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}^{A}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(c_{1}\log K)^{2}},\frac{t_{1}}{c_{1}\log K}\bigg{)}\bigg{)}.$
Setting $t_{1}=tN$, we have
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{i=1}^{N}X_{i}^{A}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
This concludes the proof for Theorem 1 part (a). ∎
A.1.4 Proof of Theorem 1 - part (b)
Similar to the proof of part (a), here we provide a high-level structure for our proof. First, recall from the discussion under Assumption 1, as proved in Equation 5, that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $-\log(\epsilon)$, and therefore it is finite. Then, In Lemma 4, we show that the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale. In Lemma 5, we find the sub-exponential norm for the martingale increments random variable $X_{n}^{A}$. As the last step, we show that martingale sum $\sum_{n=1}^{N}X_{n}^{A}$ is sub exponential with parameters $(\nu^{2}N,\nu)$, where $\nu=-c_{3}\log(\epsilon)$. Finally, we apply the tail bound for sub-exponential random variables to bound the probability of $|\sum_{n=1}^{N}X_{n}^{A}|$ being larger than a constant $t$.
Lemma 4.
Under assumption 1, the random variable $S_{N}^{A}=\sum_{i=1}^{N}X_{i}^{A}$ forms a martingale.
Proof.
$Z_{n}^{A}$ is a positive random variables, so $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is also positive. In Equation (5), we have shown that $\mathbb{E}_{p_{n}^{A}}[Z_{n}^{A}]$ is upper bounded by $-\log(\epsilon)$, and therefore we have $\mathbb{E}\big{[}|X_{n}^{A}|\big{]}\leq-\log(\epsilon)$. For a random variable to form a martingale, the following two properties need to be satisfied: (i) $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}<\infty$, and (ii) $\mathbb{E}[S_{N+1}^{A}|S_{N}^{A}]=S_{N}^{A}$.
Property (i) is satisfied because $\mathbb{E}\big{[}|S_{N+1}^{A}-S_{N}^{A}|\big{]}=\mathbb{E}\big{[}|X_{N+1}^{A}|\big{]}\leq-\log(\epsilon)<\infty$.
Property (ii) is satisfied because the martingale increments $X_{n}^{A}$ are, by definition, a zero-mean random variable conditioned on past tokens. ∎
Lemma 5.
Under Assumption 1, the sub-exponential norm for the random variable $X_{n}$ equals $-4\log(\epsilon)$.
Proof.
For the random variable to be sub-exponential, by definition 1, we need to find $t$ such that
$\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})e^{\frac{\big{|}-\log p^{A}_{n}(y_{n})-\mathbb{E}[Z_{n}^{A}]\big{|}}{t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{p^{A}_{n}(y_{n})^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]}{t}}\sum_{y_{n}\in\mathcal{X}}\frac{p^{B}_{n}(y_{n})}{\epsilon^{1/t}}\leq e^{\frac{\mathbb{E}[Z_{n}^{A}]-\log(\epsilon)}{t}}\leq 2.$
Thus, for $X_{n}$ to be sub-exponential with norm $t$, it is sufficient to satisfy
$\frac{\mathbb{E}[Z_{n}^{A}]-\log(\epsilon)}{t}\leq\log(2)\leq 1/2.$ (11)
Recall equation 5 that states
$\mathbb{E}_{p^{B}_{n}}[Z_{n}^{A}]=\sum_{y_{n}\in\mathcal{X}}-p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n}))\leq-\log(\epsilon).$
Substituting the result of (5) in (11) yields that $X_{n}^{A}$ has sub-exponential norm $\|X\|_{\psi_{1}}=-4\log(\epsilon)$. ∎
Proof of Theorem 1 part (b).
Proof.
First note that, by Lemma 5 and Lemma 1, there exists $c_{3}>0$ s.t. for $\nu=-c_{3}\log(\epsilon)$, and for any $n\in\mathbb{N}$,
$\mathbb{E}[e^{\lambda X_{n}^{A}}|\mathbf{Y}_{n-1}]\leq e^{\frac{\nu^{2}\lambda^{2}}{2}}\quad\forall|\lambda|\leq\frac{1}{\nu}.$ (12)
By Definition 2, we have $X_{n}\in SE(\nu^{2},\nu)$. Then, Following the exact same steps as in proof of Theorem 1-part (a), we conclude that $\sum_{n=1}^{N}X_{n}^{A}$ is sub-exponential with parameter $S_{N}^{A}\in SE(\nu^{2}N,\nu)$, where $\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Again, following the exact same steps we described in proof of part (a) , from (10) we have
$\mathbb{P}(|\sum_{n=1}^{N}X_{n}^{A}|\geq t_{1})\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t_{1}^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t_{1}}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Finally, setting $t_{1}=tN$, we obtain
$\mathbb{P}\bigg{(}\frac{1}{N}|\sum_{n=1}^{N}X_{n}^{A}|\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{t}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
This concludes the proof for Theorem 1 part (b). ∎
A.2 Statistical test proofs
In Section A.2.1, we present the proof for Proposition 1. In Section A.2.2, we first provide a proof overview in the first paragraph, and then provide the proof for Proposition 2.
A.2.1 Proof of Proposition 1
Proof.
Type I error occurs if a model $B_{j}\in\mathcal{B}$ generates the text string $\mathbf{Y}_{N}$, but for one model $A_{i}\in\mathcal{A}$, we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
This yields for some $A_{i}\in\mathcal{A}$, we have
$h_{N}(B_{j},A_{i})-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq h_{N}(B_{j},B_{j})-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B_{j},A_{i})-h_{N}(B_{j},B_{j})\geq\epsilon_{1}$. Hence, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\leq-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}.$
Equivalently, type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}+\bigg{|}-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}.$
So, type I error only occurs if, for at least one of the models $A_{i}\in\mathcal{A}$, we have either $|-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+|\mathcal{A}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1-part (a), we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(B_{j},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(B_{j},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}+2|\mathcal{A}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if a model $A_{i}\in\mathcal{A}$ generates the text $\bf{Y}_{N}$, but for a model $B_{j}\in\mathcal{B}$ we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
This yields that for some $B_{j}\in\mathcal{B}$, we have
$h_{N}(A_{i},B_{j})-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq h_{N}(A_{i},A_{i})-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Note that by Assumption 2, we have $h_{N}(A_{i},B_{j})-h_{N}(A_{i},A_{i})\geq\epsilon_{1}$. Hence, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\leq-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}.$
Equivalently, the type II error occurs only if at least for one of the models $B_{j}\in\mathcal{B}$ we have
$\epsilon_{1}\leq\bigg{|}-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}\bigg{|}+\bigg{|}-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}\bigg{|}.$
So, the type II error only occurs if at least for one of the models $B_{j}\in\mathcal{B}$, we have either $|-h_{N}(A_{i},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(A_{i},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type II error as
$|\mathcal{B}|\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1 we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}-h_{N}(A_{i},B_{j})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Also, from Theorem 1 part (a) we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}-h_{N}(A_{i},A_{i})(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Therefore, the type II error is at most
$2|\mathcal{B}|\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.2.2 Proof of Proposition 2
In this section, we first present Lemma 6 that provides an upper bound on the probability that, under Assumption 3, the absolute value of the difference between the cross-entropy of the (unknown) generator model $B$ and the evaluator model $A$ and the entropy of the evaluator model $A$ over string $\mathbf{Y}_{N}$ be less than or equal to a threshold $c_{4}$. In the proof of Proposition 2, we use Lemma 6 as one of the three key inequalities that yield an upper bound on type I error. Note that providing an upper bound on the type II error does not require any assumptions.
Lemma 6.
Under Assumption 3, for any positive constant $c_{4}<\epsilon_{2}$, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Proof.
We prove this lemma in three steps.
Step 1. sub-exponential norm for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
For the random variable $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$, applying (5) we obtain
$D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}=\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\big{(}\log(p^{A}_{n}(y_{n}))-\log(p^{B}_{n}(y_{n}))\big{)}\leq\sum_{y_{n}\in\mathcal{X}}-p^{A}_{n}(y_{n})\log(p^{B}_{n}(y_{n}))\leq-\log(\epsilon),$ (13)
where the first inequality holds since $\sum_{y_{n}\in\chi}p_{n}^{A}(y_{n})\log(p_{n}^{A}(y_{n}))\leq 0$, and the second inequality holds by Assumption 1.
For the random variable to be sub-exponential, by Definition 1, we need to find $t$ such that
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq 2.$
Applying (13), for the random variable to be sub-exponential with norm $t$, it is sufficient to satisfy
$\mathbb{E}\bigg{[}e^{\frac{\big{|}D_{KL}(p_{n}^{B}||p_{n}^{A})-\mathbb{E}[D_{KL}(p_{n}^{B}||p_{n}^{A})]\big{|}}{t}}\bigg{]}\leq e^{\frac{-2\log(\epsilon)}{t}}\leq 2.$
Hence, $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ has sub-exponential norm $-4\log(\epsilon)$.
Step 2. Concentration bounds for $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$
Following the same steps as in the proof for Theorem 1- part (a), we conclude that $\sum_{n=1}^{N}[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{(}-\frac{1}{2}\min\bigg{(}\frac{t^{2}}{N(-c_{3}\log\epsilon)^{2}},\frac{t}{-c_{3}\log\epsilon}\bigg{)}\bigg{)}.$
Next, setting $t=(\epsilon_{2}-c_{4})N$, we obtain
$\mathbb{P}\bigg{(}\big{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\big{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (14)
Step 3. Tail bound for cross-entropy
In this proof, for notation brevity we write $D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}$ instead of $[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$, and $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\big{]}$ instead of $\mathbb{E}\big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}\big{]}$.
First, we have
$\Big{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}$ (15)
$=$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\bigg{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-h_{N}(A,A)(\mathbf{Y}_{N})$
$\geq$ $\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}+\epsilon_{2},$
where the first equality uses the definition of cross-entropy. The first and second inequality follow from the triangle inequality. The third inequality follows from Assumption 3.
Hence, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}+\epsilon_{2}\leq c_{4}\bigg{)}$
$=$ $\mathbb{P}\bigg{(}-\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq-c_{4}+\epsilon_{2}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq\epsilon_{2}-c_{4}\bigg{)}$
$\leq$ $2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{\epsilon_{2}-c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]},$
where the first inequality follows from (15). The one-to-last inequality follows from the property of absolute value. The last inequality follows from (14). This concludes the proof. ∎
Proof of Proposition 2.
Proof.
Type I error occurs if a model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
Applying triangle inequality, we have
$\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}.$
Hence, the type I error is upper bounded as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$ (16)
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}.$
From Lemma 6, we know that for any positive $c_{4}<\epsilon_{2}$, we have
$\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (17)
From Theorem 1 we know that
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (18)
Combining Equations (16-18), we conclude the type I error is upper bounded as
$2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Setting $c_{4}=\epsilon_{2}/2$, yields that the type I error is upper bounded as
$2\exp\bigg{[}-\frac{N(\epsilon_{2}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(\epsilon_{2}/2-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}/2-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
The type II error occurs if the model $A$ generates the text $\bf{Y}_{N}$, but we have
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t.$
We upper bound the probability of this event as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$
where the inequality follows from Theorem 1-part (a). This concludes the proof. ∎
Appendix B Experiment details
B.1 Additional experiment results
AUROC Results. For this experiment, model samples are generated by sampling from the raw conditional distribution with temperature 1.
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.98* 0.84 0.99* 0.91 0.75 0.60 1.00 0.95 1.00*
Rank 0.75 0.69 0.69 0.71 0.71 0.55 0.82 0.80 0.81
LogRank 0.98* 0.87 0.99* 0.93* 0.81* 0.62 1.00 0.97* 1.00*
Entropy 0.39 0.70 0.40 0.37 0.66 0.70* 0.04 0.36 0.02
DetectGPT 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Binoculars 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98* 1.00 1.00
Diff (zero-shot) 0.01 0.04 0.01 0.06 0.18 0.27 -0.02 0.03 0.00
Roberta (base) 0.98* 0.95 0.92* 0.97* 0.92 0.69* 0.97* 0.95* 0.74*
Roberta (large) 0.98* 0.98* 0.92* 0.95 0.93* 0.68 0.96 0.93 0.65
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98 1.00 1.00
Diff (supervised) 0.01 0.01 0.08 0.02 0.06 0.28 0.01 0.05 0.26
Table 5: AUROC for detecting samples from the given models on the datasets. The best AUROC values are in bold, and the second-best values are marked with an asterisk (*). The rows Diff(zero-shot) and Diff(supervised) show our AUROC improvement over the strongest zero-shot and supervised baseline methods.
Results in Table 5 demonstrate that our method improves the average AUROC compared to the benchmarks for all datasets, with our maximum improvement obtained for the SQuAD dataset. An explanation for our higher improvement on SQuAD is the dataset’s higher entropy variance (0.128) compared to XSum (0.045) and WritingPrompts (0.074). Since our method relies on the difference between log-perplexity and the average cross-entropy, it performs more effectively when entropy’s variance is high, while aligning with likelihood-based methods when the entropy is less variable. Another observation is that our method significantly outperforms existing likelihood-based approaches in detecting Qwen-generated text for SQuAD. An explanation is that QWEN 32B is a larger model with more human-like syntax and fewer low-entropy patterns.
Convergence results. We conduct our numerical analysis for convergence on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
Same generative and evaluator model. In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (1(a)-1(d)). We consistently observe that the numerical results confirm Theorem 1- part (a) that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
(a) GPT-2 small
(b) GPT-2 medium
(c) GPT-2 large
(d) GPT-2 XL
Figure 1: Generated and evaluated by the same model
Different generative and evaluator models. To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (2(a)-2(c)). Results in these figures confirm Theorem 1. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XL
Figure 2: Generated by GPT-2 small and evaluated by different models
B.2 Dataset details
We use the codes in the official GitHub repository for DetectGPT [42], which is publicly available under the MIT license in https://github.com/eric-mitchell/detect-gpt for our white-box experiments. The datasets for this set of experiments are XSum, SQuAD, and Reddit writingPrompts that we briefly describe below. We use the codes in the official GitHub repository for the RAID dataset [11] for our black-box experiments. The RAID dataset uses a variety of datasets from 8 domains, briefly described below. For details about the RAID dataset, please refer to Appendix E.1 in [11].
XSum.
The Extreme Summarization (XSum) dataset comprises 226,711 BBC news articles paired with single-sentence summaries for each article. Each entry includes a document (the full news article), a summary (a concise, one-sentence abstraction), and an id (the BBC article identifier). The dataset is designed to support the task of extreme abstractive summarization, where models must generate highly condensed summaries that capture the essence of the source text. The XSum dataset is publicly available under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license and can be accessed at https://huggingface.co/datasets/EdinburghNLP/xsum. In our experiments, we utilize the first 30 tokens of the document column as input for creating LLM-generated text.
SQuAD.
The Stanford Question Answering Dataset (SQuAD) is a benchmark for machine reading comprehension, consisting of over 100,000 question-answer pairs derived from 536 Wikipedia articles. Each example includes a context paragraph, a question, and one or more answers, where each answer is a span of text from the corresponding context. SQuAD is publicly available under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) and can be accessed at https://huggingface.co/datasets/rajpurkar/squad. In our experiments, we utilize the first 30 tokens of the context column as input for creating LLM-generated text.
Reddit WritingPrompts.
The Reddit WritingPrompts dataset comprises over 300,000 prompt-story pairs collected from the r/WritingPrompts subreddit. Each entry includes a prompt (a creative writing prompt) and a corresponding story (a user-generated narrative response). The dataset is designed to support research in open-ended story generation and narrative modeling. The dataset is publicly available under the MIT License and can be accessed at https://huggingface.co/datasets/euclaise/writingprompts. In our experiments, we utilize the first 30 tokens of the story field as input for creating LLM-generated text.
RAID.
The Robust AI Detection (RAID) dataset is a large-scale benchmark constructed to evaluate the reliability of LLM-generated text detectors across a wide range of textual domains. It comprises over 6 million generations created by 11 language models under 4 decoding strategies and 11 adversarial attacks. RAID includes 8 carefully selected text domains to capture a broad spectrum of natural language styles, structures, and levels of creativity. The domains are
Abstracts. Sourced from arXiv.org, this domain consists of paper abstracts across multiple disciplines. Texts are technical, information-dense, and syntactically formal, reflecting real-world scientific writing in academic publishing.
Books. Drawn from the BookCorpus dataset, this domain features long-form fictional passages including dialogue and narration. It captures informal, imaginative writing with strong character and plot structure diversity.
News. Extracted from the CNN/DailyMail summarization corpus, the news domain includes professionally written news stories. These texts are structured with leads and summaries, and often exhibit a neutral tone, making them ideal for assessing generation consistency in journalistic formats.
Poetry. Taken from Project Gutenberg, the poetry domain includes English poems from public domain authors. The texts exhibit high stylistic variability, creative syntax, line breaks, and metaphorical expressions, which are characteristics that typically challenge detection.
Recipes. Based on the RecipeNLG dataset, this domain contains step-by-step cooking instructions and ingredient lists. It features a highly structured and instructional tone with imperative phrasing and domain-specific vocabulary.
Reddit. Sourced from the WritingPrompts subreddit, these texts are user-generated responses to open-ended prompts. The domain features creative fiction in a casual tone, reflecting informal narrative style and internet native conventions.
Reviews. Pulled from Amazon product reviews, this domain contains subjective, opinionated writing with informal grammar and varying tone. It reflects real-world user-generated content that combines evaluation, anecdote, and justification.
Wikipedia. This domain includes encyclopedic entries from Wikipedia, selected for their balanced, objective tone, structured format, and high factual density. Texts are varied in topic and are typically well-edited and linguistically standardized.
To avoid data contamination, all human-written documents were sourced from publicly available corpora with timestamps preceding 2022. For each document, the dataset provides one generation per model, decoding strategy, and adversarial attack. The dataset is publicly available at https://huggingface.co/datasets/liamdugan/raid under the MIT License.
B.3 The initial thirty human tokens’ impacts
Recall that for our whitebox experiments in Section 5.2, we randomly sample 500 texts from each dataset, and prompt the first 30 tokens of each text to the LLM to create an LLM-generated version. Note that the first 30 tokens in both samples by human and LLM are the same. Thus, these 30 tokens result in the same average entropy and log perplexity under both null and alternative. Denote the random variable by $C=\sum_{i=1}^{30}X_{n}$. Also, since a human generated the first 30 tokens the random variable $C$ is not zero mean. Hence, note that for both null and alternative $\sum_{n=1}^{300}X_{n}$ is biased by $C$ which is not a zero mean random variable. So, under this setting, it is not appropriate to check $|\sum_{n=1}^{300}X_{n}|$ is small to determine null. On the other hand, if we use the one-sided version of our test that rejects the null hypothesis when $\frac{1}{N}\sum_{n=1}^{N}X_{n}^{A}=\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\leq t.$ Since, the $\sum_{i=1}^{300}X_{i}$ under null and alternative is shifted by a constant $C$. The performance of the test will not be affected. Hence, we use the one-sided version of our one-sided test that removes the impact of the differences between cross-entropy and log perplexity caused by the initial 30 tokens.
B.4 Decoding strategies & Adversarial attacks
B.4.1 Decoding strategies
A decoding strategy determines how tokens are selected from the LLM’s (conditional) probability distribution in the text generation process. Greedy decoding selects the most likely (top-ranked) token from the (conditional) probability distribution. Previous research has shown that sampling from the probability distribution rather than selecting the top-ranked token (as done in greedy decoding) can make the LLM-generated text less detectable [28]. Based on these findings, [11] generates two outputs per prompt: one with greedy decoding and the other with random sampling.
The repetition penalty, as defined in [32], reduces the likelihood of generating previously used tokens through down-weighting the probability of choosing the same word in the following tokens using a parameter $\theta$. This penalty examines whether the less-repetitive output (that has undergone repetition penalty) can reduce the detectability of LLM-generated text. Formally, Given a temperature $T>0$ and a set of scores $s_{i}\in\mathbb{R}^{d}$ for each token $i$ in a vocabulary, the probability $p_{i}$ of predicting the $i^{\text{th}}$ token is
$p_{i}=\frac{\exp(s_{i}/T)}{\sum_{j}\exp(s_{j}/T)}.$
The repetition penalty defined by [32] modifies the distribution as
$\tilde{p}_{i}=\frac{\exp\left(\frac{s_{i}}{T\cdot\mathbb{I}(i\in g)}\right)}{\sum_{j}\exp\left(\frac{s_{j}}{T\cdot\mathbb{I}(j\in g)}\right)},$
where $g$ is the list of previously generated tokens, and the indicator function $\mathbb{I}(c)$ is
$\mathbb{I}(c)=\begin{cases}\theta&\text{if }c\text{ is true},\\ 1&\text{otherwise}.\end{cases}$
[11] uses [32]’s method and generates two additional outputs: with and without repetition penalty. Following [32], we use $\theta$ = 1.2 for the experiments.
B.4.2 Adversarial attacks
Adversarial attacks in LLM detection are strategically crafted paraphrases or perturbations applied to LLM-generated text that aim to bypass detection by reducing detectable differences between LLM-generated and human-written text. [11] produces different adversarially attacked versions of texts for each LLM-generated text. Below are brief descriptions of the attacks and sources from which [11] acquired them.
Insert Paragraphs [5]
[…]
Article Deletion [40, 21]
first searches through the text and captures all articles "the," "a," and "an" using POS tagging. Then, using random sampling, selects a fixed $\theta$ percent of the articles and applies the article deletion to them.
Alternative Spelling [41]
constructs a mapping between British and American spelling of words using an American to British English dictionary (https://github.com/hyperreality/American-British-English-Translator). Then, searches in the LLM-generated text to identify all the words (that also exist) in the dictionary, randomly samples a fixed $\theta$ percent of the words (with two spellings), and replaces them with the alternative spelling.
Misspelling [40, 14]
uses Wikipedia’s commonly misspelled English words dictionary (https://en.wikipedia.org/wiki/Commonly_misspelled_English_words), and finds the LLM-generated words that also exist in the misspelling dictionary.
Upper Lower Swap [14]
randomly selects a fixed $\theta$ percent of the words in the LLM-generated text, and changes the first letter to uppercase if originally lowercase, and vice versa.
Number Swap [5]
first searches through the text, and captures all numerical digits in the LLM-generated text. Then, randomly selects fixed $\theta$ percent of the digits and replace them with a randomly selected digit from 0 to 9.
Whitespace Addition [6, 14]
captures all the spaces between tokens, randomly selects a fixed $\theta$ percent of the spaces, and inserts one additional space to each selected space. Note that since the sampling is performed with replacement, the same inter-token space can be chosen multiple times, resulting in a possible insertion of more than one extra space between tokens.
Paraphrasing [35, 52]
transforms the source LLM-generated text into an alternative text with the same semantic content (i.e., the conveyed meaning is equivalent, but the content is statistically and stylistically different). [11] uses the paraphrasing method in [35]’s DIPPER-11B using their HuggingFace repository. DIPPER-11B is a paraphrasing model built on the T5-XXL architecture. The process includes two control parameters: one for lexical diversity (a measure of the variety of the words used in the text) and another for reordering, which adjusts syntactic transformation. DIPPER-11B encodes the LLM-generated text as input and then decodes a paraphrase as output. [11] uses the default setting from [35] to apply the paraphrasing attack on all texts in the dataset.
Synonym Swap [49]
does the following process (as done in [11]): first, they mask all the tokens, replacing each with a mask-fill from the top-twenty candidates in BERT [9]. Next, they assign the Part-of-Speech (POS) tag, which is a label assigned to each word in a sentence to indicate its grammatical category (such as noun, verb, adjective, or adverb), for each of the mask-fill candidates and remove all those candidates whose POS tag does not match the POS tag for the token in the input text. After creating the final list of valid candidate swaps, they select a fixed $\theta$ percent of the texts for performing the Synonym Swap adversarial attack.
Homoglyph [70, 14]
swaps some English characters with visually similar yet technically different characters from the Cyrillic scripts in [70]. Specifically, [11] uses homoglyphs for the following ASCII characters: a, A, B, e, E, c, p, K, O, P, M, H, T, X, C, y, o, x, I, i, N, and Z. They identify all these characters in the LLM-generated text, and then replace all of them (i.e., $\theta$=100%) with the homoglyph counterparts. If a character has more than one homoglyph, they randomly select between them.
Zero-width Space [21]
the Unicode zero-width space character U+200B is an invisible character that occupies space in the underlying text encoding without being visible to human readers. [11] inserts a zero-width space before and after every visible character in the perturbed text.
Note that [11]’s dataset for the Homoglyph and Zero-width space attacks is incomplete, not allowing us to provide results for them. Nevertheless, these two attacks are easy to remove with a simple pre-processing that we do in our codes (details in Appendix B.7): First, remove or normalize the invisible characters like zero-width spaces. Then map suspicious non-ASCII characters to standard ASCII or a canonical form.
To find the $\theta$ that [11] uses for each adversarial attack, please refer to Table 11 in their paper.
B.5 Detector details
RAID [11] evaluates 12 detection methods from neural classifiers, metric-based (zero-shot) detectors, and commercial detectors. All detectors were evaluated using their official implementations or web interfaces without fine-tuning on RAID data. Below is a brief description of each detector and how it was run in [11]. To check how [11] finds the thresholds to achieve FPR of 5%, please refer to Appendix F-2 in [11].
Neural classifier-based
RoBERTa (GPT2): fine-tuned to distinguish between human and GPT-2-generated text. It was evaluated using the released Hugging Face checkpoint without modification. RoBERTa-B (GPT2) is a smaller RoBERTa-base version trained on the same dataset. It is computationally lighter, but typically less accurate than the large model.
RoBERTa-B (ChatGPT) (22) fine-tuned to detect generations from ChatGPT. [11] downloads and queries the detector via HuggingFace datasets with the unique identifier Hello-SimpleAI/chatgpt-detector-roberta.
RADAR (25) a robust RoBERTa-based detector adversarially trained to resist paraphrasing and rewriting attacks. Used as released without retraining. [11] downloads and queries the detector via HuggingFace datasets with the unique identifier TrustSafeAI/RADAR-Vicuna-7B.
Metric-based (Zero-shot)
GLTR (16) evaluates the likelihood of text according to an (evaluator) LLM, bins tokens based on their likelihoods, and uses the bins as signals for detection. [11] uses the default settings from the GLTR repository (https://github.com/HendrikStrobelt/detecting-fake-text), which is setting the (evaluator) LLM to GPT2 small and the cutoff threshold at rank=10.
Fast DetectGPT (4) DetectGPT evaluates log-probability curvature under paraphrased perturbations. Low curvature regions suggest LLM generations. It was implemented using DetectGPT’s official repository with standard settings. A more efficient variant of DetectGPT that avoids expensive perturbation by approximating curvature metrics is Fast DetectGPT [4]. It was implemented using the default repository configuration, GPTNeo-2.7B as the scoring model, and GPT-J-7 B as the reference model.
Binoculars (24) uses perplexity divided by cross-perplexity for detection. RAID repository uses the code from the Binocular GitHub repository and calculates perplexity using the default models from Binocular’s repository, which are Falcon 7B and Falcon 7B Instruct.
LLMDet (71) uses proxy-perplexity, an approximation of perplexity calculated by repeatedly sampling n-grams from models rather than running the models, of the input text from 10 different small language models. None of the evaluator models used in the detector was used for text generation.
Commercial
GPTZero
([59]) relies on two metrics: perplexity, which measures how predictable the text is under an evaluator LLM, and burstiness, which captures the variation in perplexity across sentences. Human-written text typically exhibits higher burstiness, while LLM-generated text tends to be more predictable [60]. [11] queried the detector using the v2 API (https://api.gptzero.me/v2/predict/text) and threshold on the completely_generated_prob field.
Originality
[11] queried this detector through the v1 API (https://api.originality.ai/api/v1/scan/ai) and threshold on the score field in the output JSON. They queried the multilingual "version 3” of the detector for the Czech and German news domains.
Winston
[11] queried this model through the v1 API (https://api.gowinston.ai/functions/v1/predict), set the input language to English, unless the text is in German, where they set the language to German. They threshold on the score field in the output JSON.
ZeroGPT
[11] queried the API (https://api.zerogpt.com/api/detect/detectText), and used the isHuman field as the classifier output.
B.6 Hardware
All experiments were conducted using a RunPod cloud instance configured with a B200 GPU, 28 vCPUs, and 283 GB of RAM. The setup provided 150 GB of disk and pod volume and achieved bandwidth speeds of up to 20,801 Mbps in the EU-RO-1 region. Experiments were executed in an on-demand secure cloud environment available at https://www.runpod.io. Our complete set of experiments required a total of approximately 12 hours of runtime. Because our method is nonparametric and does not require training, the runtime scales linearly with the number of input texts in the dataset. Figure 3 demonstrates the RunPod cloud configuration.
Figure 3: RunPod cloud configuration used for all experiments.
B.7 Code instructions
Contents
We include our self-contained codes as part of our supplementary materials to enable replication of all experiments presented in this work. The repository comprises three Jupyter notebooks: Whiteboxattribution.ipynb for identifying the generating model among two candidate sets (codes for the experiments in Section 5.1), Whiteboxdetection.ipynb for statistical detection tests in a white-box setting (codes for the experiments in Section 5.2), and Blackbox_adversarial.ipynb for evaluating black-box detection and robustness against adversarial attacks (codes for the experiments in Section 5.3). All experiments were run on RunPod (B200) and used Hugging Face models and datasets where applicable.
How to run
Each notebook is fully self-contained and does not require external data downloads or model training, facilitating straightforward execution on personal devices. Each notebook includes
  • Installation and environment setup (in the first cells)
  • Data loading and preparation
  • Full experiment code with results
No setup beyond running the notebook cells is required since all dependencies are installed inline.
Datasets
Most datasets used in the experiments are downloaded programmatically from Hugging Face.
However, if you want to run the WritingPrompts experiments, you will need to download the WritingPrompts dataset manually from https://github.com/facebookresearch/WritingPrompts. Then, you need to place the downloaded data in the following path: data/writingPrompts/.
Portions of this code are adapted from DetectGPT, available at https://github.com/eric-mitchell/detect-gpt/tree/main. We thank the authors of DetectGPT for open-sourcing their codebase.
Appendix C Literature review
Neural classifier-based detection methods.
Neural classifier-based methods learn to differentiate between human-written and LLM-generated text through labeled examples of human-written and LLM-generated text in the training set. For example, GPT2 Detector (Solaiman et al. [56]) fine-tunes RoBERTa on the output of GPT2, and ChatGPT Detector (Guo et al. [22]) fine-tunes RoBERTa on the HC3 dataset in Guo et al. [22]. Many other supervised detection methods also exist, including those that leverage neural representations (e.g., Bakhtin et al. [3], Uchendu et al. [63]) and those that leverage bag-of-words (Fagni et al. [12]). Although some supervised detection methods show high performance, it has been observed that they tend to overfit their training domains and source models (Bakhtin et al. [3], Uchendu et al. [63]). A key concern with neural-based methods (e.g., feature-based classifiers) is their poor robustness, e.g., against ambiguous semantics (Schaaff et al. [53]), their limited ability to detect LLM-generated misinformation (Schuster et al. [54]), and their limited generalizability. Liang et al. [41] notes that since (neural-based) detectors are often evaluated on simple datasets, their performance on out-of-domain samples is typically abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using neural-based detectors. In summary, there are limitations for the neural classifier-based methods, including challenges in training models with the hype in the development of new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns. Recent efforts have been enhancing training methodology (e.g., Kumarage et al. [37], Tu et al. [62]). However, these methods provide heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts.
Metric-based/ Zero-shot detection methods.
Another stream of work attempts to distinguish machine-generated from human-written texts using statistics-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. Among those, white-box detection methods are closely related to our work. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures, e.g., as in Solaiman et al. [56]. Other measures include using GTLR based on rank-likelihood (Gehrmann et al. [16]), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. [58], entropy and Kullback-Leibler (KL) divergence (Lavergne et al. [38]), and perplexity (Vasilatos et al. [65], Wang et al. [68]). In a more recent paper, Mitchell et al. [42] observes that machine-generated texts tend to lie in the local curvature of the log probability and proposes DetectGPT, which has demonstrated reliable performance but has some limitations. Du et al. [10] analyzes the limitations of DetectGPT, highlights DetectGPT’s need for computing many perturbations, making it a computationally intensive algorithm, and proposes targeted masking strategies (rather than random masking) to improve DetectGPT’s performance. To reduce the DetectGPT’s perturbation computation cost, Bao et al. [4] introduces FastDetectGPT, which substitutes DetectGPT’s perturbation step with Conditional Probability Curvature (CPC) that accelerates the detection process by a factor of 340. Most methods focus on a single statistical property. Hans et al. [24], however, introduces a normalized measure of perplexity by dividing log-perplexity by average cross-entropy. The rationale is that LLM text is predictable to the LLM (low perplexity), whereas human text is more surprising (higher perplexity). Yet, consideration of high perplexity as a human authorship signal is unreliable due to prompt dependency [^4] This is illustrated by the ”Capybara Problem” in Hans et al. [24]. Given the prompt ”Can you write about a Capybara astrophysicist?”, an LLM may generate ”The Capybara studied dark matter in Andromeda.” Without the prompt, ”capybara” and ”astrophysicist” seem unexpectedly high in perplexity, falsely suggesting human authorship.. Similar to supervised detection methods, most statistics-based detection methods provide only heuristics.
Watermarking.
Another detection approach is to record (Krishna et al. [36]) or watermark (Kirchenbauer et al. [33]) all generated text. The method proposed in Kirchenbauer et al. [33] takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. Note that watermarking requires the cooperation of the AI company/LLM owner to implement the green listing-red listing algorithm. For a survey on watermarking methods, we refer readers to Amrit and Singh [1].
Robustness to adversarial attacks.
Growing literature examines the susceptibility of LLM detectors to adversarial attacks. Krishna et al. [35] introduces DIPPER-11B, showing that controlled paraphrasing can significantly reduce detection accuracy for different methods, e.g., DetectGPT and GPTZero. Gagiano et al. [14] benchmarks multiple detectors under a wide range of black-box attacks, including backtranslation, synonym swaps, and whitespace addition, and reports consistent detection failure across commercial and open-source detectors. Ippolito et al. [28] demonstrates that simple perturbations, such as changing punctuation or adding unrelated phrases, helps bypass classifiers like GPT-2 detectors and GLTR. Sadasivan et al. [52] proposes multiple realistic adversarial perturbations, including paraphrasing and word deletion and finds significant performance degradation of detectors under such adversarial attacks.
Possibility of detection.
As a theoretical work on the possibility of detection, Chakraborty et al. [8] provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$  [^5] Recall that in a binary binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies the model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR).. For a survey on the possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. [17].
Authorship attribution.
The literature on human-written text author attribution is extensive and, similar to the recent literature on LLM text detection and attribution, has seen developments over time, e.g., use of statistical hypothesis testing versus discriminative methods like support vector machines, neural networks, etc. In author attribution with only human authors, there is no access to the text generation "model," and the attribution is only based on the samples of texts written by different human authors. For surveys on the different methods for human author attribution, we refer readers to Juola et al. 31, Stamatatos 57 and Koppel et al. [34]. For a comparison of different methods, we refer readers to Grieve [20].
With the advent of LLMs, human author attribution extends to (1) human-written vs. machine-generated detection, (2) machine-generated vs. another machine-generated attribution, and (3) attribution in text generated through LLM and human collaboration. These shifts necessitate transitioning from human-centric stylometry to detection frameworks tailored for LLMs. For a survey that connects the older literature on authorship attribution to modern problems in the era of LLMs, we refer readers to Huang et al. [26].
Statistical tests on LLM-generated text.
Recent research has introduced statistical testing for detecting LLM-generated text and verifying whether the output behavior of deployed (e.g., cloud-hosted) models remains consistent with a known reference model. Li et al. [39], focuses on detecting watermarked text. Their method uses the next-token probabilities (NPTs) to construct likelihood-based hypothesis tests. Gao et al. [15] determines whether a commercial or third-party API provider has modified the underlying language mode, which is a practical concern when users rely on cloud-accessed LLMs whose internals may be updated without notice. They formalize this as a two-sample hypothesis testing problem, where users compare outputs from the deployed API against samples from a previously archived reference model. To quantify distributional shifts, they propose using Maximum Mean Discrepancy (MMD) with a string kernel, a nonparametric method first introduced in Gretton et al. [19], which enables detection of subtle changes in the model’s token distribution. Our framework is zero-shot, and as a result, it requires neither labeled LLM outputs nor access to the generation model (as elaborated in Section 4.2), and we also theoretically show that the type I and type II errors of our test decrease exponentially in the text length.
Asymptotic features of LLM-generated text.
Mudireddy et al. [43] shows that the log-perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions when the evaluator and the generator model are the same. Their result relies on sufficiently large text samples to reach asymptotic behavior. This version of their manuscript, with correct proofs, was made online at the same time/after the initial release of our manuscript. Our work is a generalization to the case with an arbitrary length text and the case where generator and evaluator models are different by showing the convergence of log-perplexity to cross-entropy between the generator and the evaluator models, except with an exponentially small probability in the string length.